### v1

# PixFoundation: Are We Heading in the Right Direction with Pixel-level Vision Foundation Models?

Abstract
Multiple works have emerged to push the boundaries on multi-modal large language models (MLLMs) towards pixel-level understanding. Such approaches have shown strong performance on benchmarks for referring expression segmentation and grounded conversation generation. The current trend in pixel-level MLLMs is to train with pixel-level grounding supervision on large-scale labelled data. However, we show that such MLLMs when evaluated on recent challenging vision centric benchmarks, exhibit a weak ability in visual question answering. Surprisingly, some of these methods even downgrade the grounding ability of MLLMs that were never trained with such supervision. In this work, we propose two novel challenging benchmarks and show that MLLMs without pixel-level grounding supervision can outperform the state of the art in such tasks when evaluating both the pixel-level grounding and visual question answering. We propose simple baselines to extract the grounding information that can be plugged into any MLLM, which we call as PixFoundation. More importantly, we study the research question of “When does grounding emerge in MLLMs that are not trained with pixel-level grounding supervision?” We show that grounding can coincide with object parts or location/appearance information. Code repository is at https://github.com/MSiam/PixFoundation/.
1 Introduction
There have been numerous advancements in pixel-level image and video understanding, including tasks such as image/video segmentation (Zhou et al., 2022; Minaee et al., 2021; Kirillov et al., 2023; Ravi et al., 2024), pixel-level visual grounding and reasoning (Rasheed et al., 2024; Lai et al., 2024), depth estimation (Yang et al., 2024) and tracking (Wang et al., 2023). The majority of these have been transformed with the emergence of foundation models (Bommasani et al., 2021), specifically multi-modal large language models (MLLMs) (Liu et al., 2023/; Dai et al., 2023). Nonetheless, pixel-level MLLMs have shown degradation in their capabilities in chat performance (Lai et al., 2024). Recent models tried to address this gap (Zhang et al., 2024b, a), yet they relied on standard evaluation benchmarks overlooking the shortcomings of current MLLMs.
Recent efforts explored the shortcomings of MLLMs in vision centric benchmarks (Tong et al., 2024b, a). Such benchmarks focus on challenging visual patterns and tasks such as counting and relative positioning. Nonetheless, these benchmarks did not evaluate the recent pixel-level MLLMs. In this work we aim to provide challenging vision centric benchmarks that are dedicated to evaluate these models. Through these, we answer the first research question; “Are the current pixel-level MLLMs trained with full grounding supervision heading in the right direction to improve both grounding and visual question answering (VQA) among other capabilities?”. Our findings show that the majority of pixel-level MLLMs still fall short in such challenging setting. While evidently, some of these show superior performance in visual grounding, we show that MLLMs that were not trained with pixel-level grounding can have better performance without degrading their capabilities.
There have been recent works showing training-free open vocabulary segmentation emerging from vision language models (Wang et al., 2024; Luo et al., 2024; Hajimiri et al., 2025). Concurrent work has specifically explored emerging grounding in MLLMs (Cao et al., 2024). We are inspired by the previous work to establish a baseline on our proposed benchmarks. However, unlike their method we focus on the second research question of “When does grounding emerge in MLLMs that are not trained with full grounding supervision?”. Previous efforts studying emergent segmentation and/or grounding assumed correspondence with the specific language tokens (i.e., nouns or noun phrases) of the objects of interest. However, our work documents that emerging grounding in MLLMs does not necessarily coincide with the exact language tokens of the object. We show it mostly emerges in the last 40% of the output, and can coincide with concepts about object parts, position/location, color/appearance or context of these objects. Fig. 1 summarizes our two major research questions.
In summary, our contributions include: (i) Providing pixel-level segmentation annotations and referring expression of the object of interest in the question for two vision centric benchmarks (Tong et al., 2024a, b), which are PixMMVP and PixCV-Bench. (ii) Benchmarking recent efforts in pixel-level MLLMs where we show that they degrade VQA capabilities. More importantly, some of them lag in visual grounding with respect to simple techniques of extracting the segmentation from vanilla MLLMs, i.e., MLLMs that are not trained for pixel-level grounding. (iii) We provide a simple mechanism for extracting segmentation from vanilla MLLMs, with an understanding of when grounding emerges. Our mechanism uses the observation that grounding can emerge corresponding to different output tokens describing the object’s appearance or location, not necessarily the exact text of the object of interest. We call it PixFoundation, since it mines the foundation model for pixel-level understanding.
2 Related Work
Pixel-level Vision Foundation Models. There have been various vision foundation models released that were trained with supervision for the segmentation task (e.g., SAM, SAM 2.0) (Kirillov et al., 2023; Ravi et al., 2024). Orthogonal to this, some methods discussed the ability of vision foundation models such as CLIP and BLIP in image segmentation without any segmentation supervision (Luo et al., 2024; Hajimiri et al., 2025; Wang et al., 2024). Yet, they relied on earlier foundation models that did not incorporate the power of large language models. Combining large language models with vision has been extensively researched with pioneering works such as Llava (Liu et al., 2023/, 2024) and instruct-BLIP (Dai et al., 2023). Multiple works afterwards focused on pixel-level visual grounding in these MLLMs with full supervision (Lai et al., 2024; Rasheed et al., 2024; Zhang et al., 2024a, b, a, b). However, these methods were lagging in their chat performance. Notably, pixel-level MLLMs were not evaluated on the challenging benchmarks that focused on the shortcomings of MLLMs (Tong et al., 2024b, a). Hence, it is still unclear if the pixel-level grounding supervision helped to improve their ability on these challenging tasks or not. In this work, we focus on the previous question to have a better understanding of their performance. Concurrent work, has shown that without pixel-level supervision there is an emerging ability to perform pixel-level grounding (Cao et al., 2024). We rely on this method as our baseline, but unlike previous works we provide an insight on when grounding emerges in such MLLMs.
Benchmarking Multi-modal Large Language Models There is an abundance of standard benchmarks used for evaluating MLLMs (e.g., MMU (Yue et al., 2024)) and their pixel-level counter parts (e.g., refCOCO/+/g (Yu et al., 2016; Kazemzadeh et al., 2014)). These benchmarks have pushed the limits on MLLMs capabilities in terms of VQA and visual grounding. Nonetheless, there have been various works that discussed the shortcomings of MLLMs. In one of these works they discussed specifically CLIP (Radford et al., 2021), which is used in various MLLMs as the visual backbone. They proposed a benchmark, MMVP (Tong et al., 2024b), that is focused on the visual aspects within a VQA task designed on CLIP blind pairs. More recently, CV-Bench (Tong et al., 2024a) focused on two major tasks that are vision focused which are counting and relative positioning. Both were proposed to evaluate MLLMs that do not have the ability to generate segmentation output. Nonetheless, they still provide quite challenging scenarios that can act as a strong benchmark for the pixel-level MLLMs counterpart. In this work, we extend these two benchmarks to augment it with pixel-level annotations and referring expressions that correspond to the object of interest within their VQA task. These help to understand if failures occurring in these benchmarks stem from grounding or other reasons.
3 Method and Benchmarks
In this section, we describe our two benchmarks and probing techniques for pixel-level MLLMs and MLLMs that were not trained with pixel-level grounding supervision.
3.1 Benchmarks
PixMMVP benchmark: We build upon the recently released MMVP (Tong et al., 2024b) which identified clip blind pairs and used them to build a challenging benchmark with the corresponding questions and choices for 300 images. We augment the aforementioned dataset and manually annotate each question with the corresponding object of interest referring expression, e.g. an elderly person or the butterfly’s feet. There are seven questions only that are not designed to inquire about a specific object in the scene, which are excluded. Examples include questions inquiring on the view direction of the camera which are not tied to a specific entity. Our manual referring expression annotations are as fine-grained as possible. These expressions correspond to what needs to be grounded in the image to answer the question. Afterwards, we manually label these objects of interest with polygonal annotations using the VGG annotator (Dutta et al., 2016).
PixCV-Bench benchmark: For this benchmark we build upon the 2D component of the recently released CV-Bench (Tong et al., 2024a). We specifically select the 2D component, since they are sourced from segmentation datasets (i.e., ADE20K (Zhou et al., 2017) and COCO (Lin et al., 2014)), which can be used in our proposed benchmark. However, the publicly released CV-Bench does not identify the objects in question and their corresponding segmentation. As such we use GPT-4o to parse the questions and identify the objects of interest automatically, followed by manual inspection and correction. Specifically, we collect the classes in each image from the corresponding dataset and construct a list of class choices “1. CLS1, 2. CLS2, …”. Then we prompt GPT-4o with the following, “Provide number only as an answer. Identify the objects of interest in the following question: QUESTION ? 1. CLS1, 2. CLS2, … ”. This provides us with the categories per question that highlights the objects of interest. While seemingly these are categorical annotations not referring expressions, certain scenarios in CV-Bench are different. Specifically, in the relative positioning task all the questions that include an object highlighted by a red box in the image are annotated with the referring expression, “(annotated by the red box)”, beyond simple categorical annotations.
Afterwards, we use the selected categories from GPT-4o to retrieve the corresponding segmentation mask/s per image. Furthermore, we use a custom annotation tool to manually filter the objects in the question, e.g. selecting only the object mask annotated by the red box when referred to it and filtering out the other instances for that same class. Another example that needs manual filtration when the class in question is a broader category than what is inquired upon, e.g., “Pendant Lamp” which is under the category of “Lamp” in ADE20K. In such a case we need to filter out the masks of other types such as “Table Lamp”. Moreover, we identified missing annotations in rare occasions that required additional intervention. Examples include transparent objects such as wine glasses which were partially occluded, yet the ground-truth answer of the CV-Bench question included that object in the count. As such, we use VGG annotator for these missing annotations that were not part of the original segmentation masks. We provide the final PixCV-Bench with referring expressions and their segmentation annotations that can be used to evaluate the grounding ability in relation to the original VQA task. Appendix A provides visual examples from our benchmarks.
3.2 A Pixel-level MLLMs Study
We utilize the two proposed benchmarks, PixMMVP and PixCV-Bench, to evaluate how the current trend in pixel-level MLLMs that relies on training with grounding supervision perform in such challenging tasks. Furthermore, we inspect the failures of these pixel-level MLLMs and explore simple approaches to pixel-level understanding from MLLMs that overcome the previous shortcomings.
Pixel-level MLLMs Shortcomings. We highlight the failures for the current state-of-the-art pixel-level MLLMs through three probing techniques. First, we highlight the degraded performance in VQA from most of the pixel-level MLLMs that are trained with pixel-level grounding supervision. We use for the first prompt “IMGQUESTION? OPTION1 OPTION2…”, as shown in Figure 2. Certain pixel-level MLLMs tend to answer the aforementioned question while outputting a corresponding segmentation mask/s for the objects of interest. Notably, the worst two models in this task, LISA (Lai et al., 2024) and GLAMM (Rasheed et al., 2024), are not able to provide an answer and rather refer to a segmentation mask. On the other hand, OMG-Llava (Zhang et al., 2024b) shows better ability in visual question answering.
The second shortcoming we discuss is their degraded ability to visually ground objects. Surprisingly, although they were trained with pixel-level grounding supervision, not all of these models show superior grounding performance. Figure 3 shows the second prompt to generate a segmentation mask for the ground-truth referring expression. The purpose of this probing is to understand whether the failure in these models is purely on the VQA task, or its inability to ground the objects of interest in the corresponding question or both. Figure 3 shows the worst two models in this aspect, which are GLAMM, the region captioning variant, and Llava-G. Both fail to segment the specific object in question, while OMG-Llava shows better performance in grounding.
Third, we highlight another shortcoming, where these MLLMs exhibit degraded ability to follow instructions. In order to probe this, we use the following prompt: “IMGQUESTION? a.OPTION1 b.OPTION2… Answer with the option’s letter from the given.” Figure 4 shows an example with the answers from the worst two models in this aspect which are LISA (Lai et al., 2024) and Llava-G (Zhang et al., 2024a). Both are incapable of following the instruction, yet Llava-G tries to tackle the question unlike LISA. On the other hand, OMG-Llava shows better ability to follow the instruction and answer the question.
Baselines and upper bounds. In addition to evaluating state-of-the-art pixel-level MLLMs, we propose two baselines and one upper bound. The first of which is inspired by a concurrent work (Cao et al., 2024) that identified the emergent grounding in multi-modal large language models without the need for any pixel-level grounding supervision. Specifically, we use their attend and segment meta architecture as one of our baselines. However, we are the first to discuss when does such grounding emerge in these models. We identify an interesting connection between the identified output tokens and the output grounding from the attention maps that gives insights on how these models reason.
The attend and segment meta-architecture extracts the raw attention map for the output token, , where are the number of layers and heads, resp. Then, are the number of input language tokens before and after the visual tokens respectively, while are the height and width of the input image. Only the attention corresponding to the visual tokens of length are used, and these attention maps are averaged across the layers and heads, resulting in . This is further normalized across all the output, for output tokens. The attend and segment depends on using spaCy natural language processing tool (Honnibal et al., 2020) to identify the noun phrases and associate them to the ground-truth referring expressions. Thus, the spaCy embeddings closest to the ground-truth expression are used in the mask selection. This is followed by extracting the maximum attention point to feed into SAM (Kirillov et al., 2023) as a point prompt.
| Method | PixGr. Train | MMVP & PixMMVP | ||||
| Llava 1.5 (7B) (Liu et al., 2023/) | ✗ | 28.7 | 28.0 | - | - | - |
| Llava 1.5 (13B) (Liu et al., 2023/) | ✗ | 39.3 | 30.0 | - | - | - |
| Cambrian (8B)* (Tong et al., 2024a) | ✗ | 52.0 | 52.0 | - | - | - |
| OMG Llava (7B) (Zhang et al., 2024b) | ✓ | 12.0 | 12.0 | 17.8 | 38.0 | 18.2 |
| GLAMM (7B) (Rasheed et al., 2024) | ✓ | 1.3 | 2.7 | 31.5 | 47.4 | 5.1 |
| GLAMM - RegCap (7B) (Rasheed et al., 2024) | ✓ | 12.7 | 6.7 | 14.5 | 18.6 | 15.1 |
| LISA (7B) (Lai et al., 2024) | ✓ | 7.3 | - | 18.1 | 42.9 | 12.5 |
| Llava-G (7B) (Zhang et al., 2024a) | ✓ | 9.3 | - | 17.8 | 13.5 | 12.2 |
| Llava 1.5 (7B) + (a+s) (Cao et al., 2024) | ✗ | 28.7 | 28.0 | 11.1 | 11.2 | 16.1 |
| Llava 1.5 (13B) + (a+s) (Cao et al., 2024) | ✗ | 39.3 | 30.0 | 9.8 | 11.4 | 17.7 |
| Cambrian (8B)* + (a+s) (Cao et al., 2024) | ✗ | 52.0 | 52.0 | 14.3 | 15.1 | 23.4 |
| PixFoundation (Llava (7B)) (Ours) | ✗ | 28.7 | 28.0 | 16.9 | 18.8 | 22.7 |
| PixFoundation (Llava (13B)) (Ours) | ✗ | 39.3 | 30.0 | 14.4 | 18.2 | 24.9 |
| PixFoundation (Cambrian (8B)*) (Ours) | ✗ | 52.0 | 52.0 | 17.2 | 18.9 | 27.7 |
| Upper Bound - Oracle Selection | ||||||
| PixFoundation (Llava (7B)) (Ours) | ✗ | 28.7 | 28.0 | 26.1 | 38.0 | 32.7 |
| PixFoundation (Llava (13B)) (Ours) | ✗ | 39.3 | 30.0 | 23.6 | 38.2 | 38.7 |
| PixFoundation (Cambrian (8B)*) (Ours) | ✗ | 52.0 | 52.0 | 52.0 | 56.1 | 54.0 |
For our baseline and upper bound, we build upon the previous pipeline and build an oracle upper bound and an automatic baseline. We introduce two main modifications to account for our observation that the correct grounding can occur with different output tokens describing the object not necessarily aligning with the exact ground-truth expression. The first modification is to go through all the potential output tokens without relying on spaCy embeddings. In the oracle we rely on the ground-truth mask to select the correct token and its corresponding attention map with highest intersection over union as an upper bound. The automatic baseline uses a simple but powerful mechanism where we overlay the predicted masks on the original image to highlight the potential object of interest. This is followed by feeding these images to a multi-modal LLM inquiring on which is best in highlighting this object. Specifically, we use the following prompt “Select the image that has EXPR best highlighted in red color than the others? Answer with a number from 1 to N and mention the number only. IMG”, where EXPR and IMG are the ground-truth expression and the image tokens respectively. In our experiments we rely on the open source Cambrian-1(8B) (Tong et al., 2024a) model in our automatic baseline for the mask selection. The second modification since SAM has a good understanding of point prompting ambiguity, we process three potential output masks for each prompt instead of a single one. This enables us to utilize the power of SAM in identifying fine-grained objects and referring expressions that tends to surpass what other MLLMs do, even those trained with pixel-level grounding supervision.
| Method | PixGr. Train | CV-Bench & PixCV-Bench | ||||
| Llava 1.5 (7B) (Liu et al., 2023/) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | - | - | - |
| Llava 1.5 (13B) (Liu et al., 2023/) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | - | - | - |
| Cambrian (8B)* (Tong et al., 2024a) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | - | - | - |
| OMG Llava (7B) (Zhang et al., 2024b) | ✓ | 9.2/14.7/12.0 | 36.8/47.4/42.1 | - | 50.5 | 45.9 |
| GLAMM (7B) (Rasheed et al., 2024) | ✓ | - | - | 30.2 | 51.9 | - |
| GLAMM - RCap (7B) (Rasheed et al., 2024) | ✓ | 22.8/32.8/27.8 | 46.8/62.0/54.4 | 3.6 | 7.4 | 13.0 |
| LISA (7B) (Lai et al., 2024) | ✓ | 1.9/5.5/3.7 | - | 16.8 | 48.1 | 6.7 |
| Llava-G (7B) (Zhang et al., 2024a) | ✓ | 13.9/14.2/14.1 | 5.1/3.7/4.4 | 1.7 | 17.6 | 15.8 |
| Llava 1.5 (7B) + (a+s) (Cao et al., 2024) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | 5.2 | 15.6 | 24.8 |
| Llava 1.5 (13B) + (a+s) (Cao et al., 2024) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | 4.7 | 14.9 | 24.0 |
| Cambrian (8B)* + (a+s) (Cao et al., 2024) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | 18.6 | 15.9 | 29.6 |
| PixFoundation (Llava (7B)) (Ours) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | 5.3 | 19.1 | 29.0 |
| PixFoundation (Llava (13B)) (Ours) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | 4.7 | 17.7 | 27.5 |
| PixFoundation (Cambrian (8B)*)(Ours) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | 12.1 | 16.6 | 27.0 |
| Upper Bound - Oracle Selection | ||||||
| PixFoundation (Llava (7B)) (Ours) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | 6.3 | 49.7 | 54.5 |
| PixFoundation (Llava (13B)) (Ours) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | 5.3 | 50.6 | 55.4 |
| PixFoundation (Cambrian (8B)*) (Ours) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | 54.3 | 64.4 | 68.1 |
the dorsal fin of the animal
4 Experiments
4.1 Experimental Setup
Evaluation benchmarks, protocols and metrics. PixMMVP is composed of 300 images paired with questions, choices, referring expressions and segmentation masks, while PixCV-Bench has 1,438 images with their corresponding annotations similarly. On each benchmark we evaluate the VQA and visual grounding capabilities following three probing techniques and reporting their metrics. The first probing is to evaluate the VQA ability, where the accuracy is computed using GPT-4o following (Tong et al., 2024b) as, . If the model generates a segmentation without explicitly asking it to, it is evaluated with respect to the ground-truth referring segmentation in terms of mean intersection over union as, . The second probing prompts the model to identify the referred expression then evaluates the mean intersection over union reported as, . The third probing following (Tong et al., 2024a) instructs the model to generate a single option letter and evaluate the accuracy directly without GPT-4o, reported as, . There is a need for the first probing since some of the recent pixel-level MLLMs face challenges in following instructions. We evaluate the score of each model, , which is the harmonic mean across the maximum of both pixel-level visual grounding and VQA,
We mainly focus on evaluating four state-of-the-art pixel-level MLLMs; LISA (Lai et al., 2024), GLAMM (Rasheed et al., 2024), OMG-Llava (Zhang et al., 2024b) and Llava-G (Zhang et al., 2024a). For GLAMM we use two variants; the original model (GLAMM) and the one fine-tuned for region-level captioning, (GLAMM-RegCap). We use Hugging Face for these models weights, refer to Appendix A.
Baselines and upper bound implementation details. We evaluate: (i) the attend and segment (a+s), (ii) the oracle selection relying on the highest intersection over union in selecting the correctly predicted masks (PixFoundation), and (iii) the automatic selection, (PixFoundation). These are implemented on top of three base MLLMs which are, Llava 1.5 (7B, 13B) (Liu et al., 2024) and Cambrian-1(8B) (Tong et al., 2024a). The automatic selection is implemented using the Cambrian-1 (8B) model, yet it can be further improved using GPT-4o, refer to Appendix A for more details.
4.2 Are the current pixel-level MLLMs heading in the right direction?
In order to answer this, we evaluate each of these pixel-level MLLMs capability in VQA in challenging tasks. Additionally, we evaluate their ability to visually ground the objects of interest in these questions.
PixMMVP Benchmark. Table 1 shows the results on the challenging PixMMVP benchmark. From the accuracy of VQA, MLLMs that are not trained with pixel-level grounding surpass their pixel-level counterpart with up to 16%. The best in pixel-level MLLMs score in this aspect is GLAMM-RegCap (Zhang et al., 2024b) yet it has degraded ability to generate segmentation. On the other hand, when looking at pixel-level visual grounding we find the best model, GLAMM (Rasheed et al., 2024), has a weak ability in VQA or following instructions. Moreover, it shows LISA and Llava-G are incapable of following the instruction to output the option letter reported in, . Looking at the bottom three rows, the oracle confirms that MLLMs that were never trained with pixel-level grounding have the correct grounding within their learned attention maps, refer to Fig. 5. Additional qualitative analysis is in Appendix B. Looking at the final score, , the oracle variant, PixFoundation (7B), outperforms the corresponding best pixel-level MLLMs, OMG-Llava (7B), by a considerable margin while the automatic outperforms it with up to 4%. While the attend and segment baseline (Cao et al., 2024) lags behind our automatic method with more than 6%.
PixCV-Bench Benchmark. Table 2 shows the results on PixCV-Bench. Persistently all the pixel-level MLLMs still lag behind the MLLMs that were never trained with pixel grounding supervision in VQA. Nonetheless, OMG-Llava strikes the right balance in both VQA and pixel-level grounding with the highest score, , within the 7B models. While, GLAMM, shows the best pixel-level grounding yet it is incapable of VQA on this benchmark. Looking at the bottom three rows for the oracle similar findings to PixMMVP emerge. The attend and segment baseline and the automatic PixFoundation baseline, both lag behind the oracle as expected yet they show competitive performance in pixel-level visual grounding, even better than some of the MLLMs trained with pixel-level supervision. Our automatic PixFoundation outperforms the attend and segment using the Llava variants, and is competitive in Cambrian-1. Cambrian-1 is more challenging as it generates more tokens, thus selecting among higher number of masks, refer to Appendix C.
Summary. In summary, pixel-level grounding supervision degrades MLLMs ability in VQA and sometimes even their generalization in grounding. Failure case analysis is provided in Appendix D. We show that MLLMs trained with pixel-level supervision lag behind vanilla MLLMs using simple mechanisms to extract grounding, and the oracle indicates there is an opportunity to improve this. Moreover, we show that grounding might not coincide with the noun phrase most similar to the referred expression, where our oracle upper bound is surpassing the attend and segment.
4.3 When does grounding emerge in MLLMs?
When - location. Taking into account the powerful performance of the oracle upper bound, it begs the important question of when grounding emerges. We start by looking at when it emerges in terms of the location within the output text. We analyze the word/phrase location with respect to the full output text in terms of a percentage from its total length, (i.e., 0% means the beginning of the text). Accordingly, Fig. 6(a) shows the location percentages histogram, binned at 10%, for the three base MLLMs reporting the oracle selection and evaluating on PixMMVP benchmark using the second probing. In the Llava 1.5 variants, grounding mostly emerges at the last 40% of the output text, while for Cambrian it is mostly at the last 10%.
When - concept. For the second analysis we look into the concept category that the correct output word/phrase corresponds to. The previous assumption in other works is that grounding emerges in exact noun or noun phrase of the object of interest. Except our analysis confirms that this is not necessarily the case. We take the correct noun/noun phrase where the grounding emerges based on the oracle from all the three variants, then we pass it to GPT-4o to request a grouping of these concepts. Specifically, we prompt it with the following “Can you group these noun phrases into categories of abstract concepts. You are provided with pairs of (noun phrase, its full context).”, where the entire output text is the full context. After manual filtration and merging it resulted into six main groups, which are: (i) Color and appearance, (ii) location and position, (iii) object parts, (iv) context and setting, (v) objects and entities, and (vi) State. We then prompt for each of the noun/noun phrase, GPT-4o, to categorize it within these six categories. The histogram of the occurrences of these concept categories is shown in Fig. 6(b). It clearly conveys that in certain scenarios the correct output when grounding emerges can be describing the position or the location of the object of interest not necessarily the exact ground-truth referring expression. While the majority occurs under “Objects and Entities”, there are still percentage of output coinciding with position, color, context, object part, or even the state. Further qualitative results are provided in the Appendix B.
Random vs. best. All the results of our baselines and our findings hinge on the fact that we are using the maximum attention per output noun phrase to prompt SAM for the segmentation mask. Nonetheless, as a lower bound analysis, we evaluate the performance if we use a random point as prompt to SAM instead. For fair comparison, we generate random points with the count of output masks that the oracle has to select among (i.e., the number of the output noun phrases). We conduct this ablation on PixMMVP using Llava 1.5 (7B) base MLLM, with random point prompts followed by the oracle selection among their SAM masks. Figure 7 shows that this random + oracle baseline is a strong baseline, yet it lags behind the correct one using the maximum point (i.e., First) with around 12%. More importantly, we confirm the stability of the results if we select the second best or third best attention (i.e., Second and Third), which are on-par to the maximum point. Thus, even with the oracle selection, using the wrong point prompts lags with a considerable margin behind using the correct ones.
Summary. In summary, we found that grounding emerges in the last 40% of the output text which might indicate that the reasoning of the model to provide a response impacts when the correct visual grounding can occur. More importantly, we show that grounding in MLLMs can emerge in the noun phrase that corresponds to color, position or other characteristics of the object of interest and not necessarily the exact referring expression.
5 Conclusion
We proposed two benchmarks showing that pixel-level MLLMs degraded the ability in VQA and even grounding of fine-grained objects. Thus, our results questioned whether we are heading in the right direction with these models. Additionally, we provide powerful baselines with improved scores without training for pixel-level grounding.
Impact Statement
Multi-modal large language models are widely used in various applications, such as robotics, medical image processing and remote sensing. The pixel-level understanding within such MLLMs is necessary for such applications that require the localization and even in certain scenarios the delineation of the boundaries for the objects of interest. It is even more important to maintain a good chat performance and visual question answering ability in such applications as well. In our work, we have investigated the shortcomings of pixel-level MLLMs while providing more challenging benchmarks for these, to improve them further.
However, as with many other AI advancements there are risks that could be entailed from the deployment of such models. There could be inherent biases emerging in such pixel-level MLLMs impacting various under-represented groups. We think that our benchmarking efforts and providing a tool to understand the pitfalls in the understanding and reasoning of these models could be an initial direction for mitigating such biases. Nonetheless, we leave it for future work to explore this further.
Appendix A Additional implementation details
In this section, we cover additional details about our proposed datasets and the implementation of the evaluation setup and baselines. We also refer to the output from the questions of the three probing techniques in the supplementary material for all the studied models.
Datasets. Our proposed datasets, PixMMVP and PixCV-Bench, are composed of ground-truth referring expressions describing the object of interest in the respective question and its segmentation mask. We show in Fig. 8 examples of these ground-truth annotations for both datasets. It shows the challenging scenarios in pixel-level visual grounding which is strongly tied to the visual question answering task, since an integral part of answering these questions requires the grounding of the object/s of interest.
Models. We also detail the model checkpoints we use for the four pixel-level MLLMs and their variants, retrieved from HuggingFace (Wolf et al., 2019) in Table 3. These also include the model checkpoints used for the base MLLMs that were not trained with pixel-level visual grounding. It is worth noting, that for GLAMM we use two variants (FullScope and RegCap) since their base model (i.e., FullScope) has low performance in the visual question answering task. As such we use the other variant for GLAMM that was fine-tuned for region-level captioning using RefCOCOg and Visual Genome dataset. Furthermore, we provide details on the oracle selection mechanism, where in the case of multiple masks exhibiting the same IoU we select the last occurring one in the when analysis, since we need to select the location of when it occurs. While in the quantitative and qualitative evaluation, we resort to simply not selecting any mask. These occur in few cases when the ground-truth mask is all background, thus any mask IoU will result in zero.
Additionally, we provide details on the SAM model that is used in the three baselines and upper bounds in our benchmarks, where we use the ViT-H variant. Finally, we provide an illustrative example of our automatic selection mechanism with the corresponding predictions on PixMMVP using Llava 1.5 (7B) in Fig. 9. Our automatic selection goes through an iterative process of prompting the selected MLLM, in our case Cambrian-1, with three images overlayed with predicted segmentation to select the best within each group of three. In the final stage the best images are used to prompt the MLLM to select the final selected mask that best describes the object of interest. In the oracle upper bound whenever the model is to be evaluated in a multiple object scenario we take all the possible pairs of the masks then select the best pair based on the highest intersection over union.
| Model Name | Model Checkpoint |
| LISA | xinlai/LISA-7B-v1-explanatory |
| GLAMM | MBZUAI/GLaMM-FullScope |
| GLAMM-RegCap | MBZUAI/GLaMM-RegCap-RefCOCOg |
| Llava-G | Haozhangcx/llava_grounding_gd_vp |
| Llava 1.5 (7B) | 4bit/llava-v1.5-7b |
| Llava 1.5 (13B) | 4bit/llava-v1.5-13b-3GB |
| Cambrian-1 (8B) | nyu-visionx/cambrian-8b |
Evaluation. We also provide the details on computing the visual question answering accuracy using GPT-4o in the first protocol (Tong et al., 2024b). We use the following prompt: “Given the following question QUESTION, the correct answer is ANSWER. Does the following answer correctly answers the question, answer: RESPONSE? Respond with a Yes/No”.
Appendix B Additional qualitative analysis
In this section, we provide a qualitative ablation of our baselines and a visualization of the attention maps that can show how vanilla MLLMs are reasoning on the question they are answering. Additionally, we provide qualitative examples showing when grounding emerges in these vanilla MLLMs. Finally, we provide more examples on PixMMVP and PixCV-Bench benchmarks.
B.1 Baselines ablation
We show the qualitative ablation among the two baselines and upper bound using the best base MLLM Cambrian-1 (8B) in Fig. 10 on PixMMVP. The three confirm that there is grounding emerging in MLLMs that were not trained with pixel-level grounding supervision. Nonetheless, it shows that identifying when that grounding emerges is equally important in retrieving the best segmentation of the referring expression. The first baseline, attend and segment, assumes the alignment between the attention map that can be mined for the segmentation mask and the noun phrase that has the highest correspondence to the ground-truth category or noun phrase. Our findings quantitatively and qualitatively show otherwise, where grounding can emerge in different output tokens. It also shows the oracle upper bound for mask selection, PixFoundation, exhibiting better segmentation than the attend and segment, confirming on the aforementioned finding. Additionally, it shows that our simple automatic mechanism, PixFoundation, surpasses the attend and segment as well on PixMMVP.
Attend and Segment (Cao et al., 2024) (Cambrian-1 (8B))
PixFoundation (Cambrian-1 (8B))
PixFoundation (Cambrian-1 (8B))
[5pt]
the yellow skin
[5pt]
blue overalls
[5pt]
goggles
[5pt]
six legs
[5pt]
the front legs
[5pt]
a hairy texture
B.2 Attention maps visualization
In this section, we visualize the normalized attention maps, , in Fig. 11. We show two examples for Cambrian-1 (8B) from PixMMVP using the first probing where we directly prompt the model with question and options. The first row shows outstanding ability to visually ground the different noun phrases from the output text. The full output text of the first row example is: “The image provided is a cake designed to resemble a minion from the D̈espicable Mef̈ranchise. It is not a living creature and therefore cannot smile or have a tongue out. The cake is an inanimate object, crafted to mimic the appearance of a minion, which is a fictional character from the animated movie series. The design elements such as the yellow skin, blue overalls, and goggles are characteristic of the minions’ appearance in the films.” The visualization shows how the maximally attended locations for the last three noun phrases correspond to the correct locations in the image.
The second output text corresponding to the example shown is; “The butterfly’s feet, also known as tarsi, are not distinctly visible in this image due to the angle and the butterfly’s wings being open. However, we can infer their presence and approximate location. Monarch butterflies have six legs, with the hind legs being the longest and the front legs being the shortest. The legs are typically slender and have a hairy texture, which aids in gripping onto surfaces. In this image, the legs are likely located at the bottom of the butterfly’s body, just below the abdomen, and are probably in contact with the leaf it is perched on.” The attention maps highlight what we suspect is a failure where the MLLM mistakes the antenna of the butterfly for front legs. Such hidden failures that do not necessarily affect the correctness of the answer, are still important to study and we believe our tool with the oracle upper bound can be used to inspect this further. Finally, we find that these attention maps in both examples are not sufficiently accurate to be used for segmentation directly, yet when paired with a powerful segmentation method like SAM it provides a good segmentation performance.
B.3 When does grounding emerge?
We show additional examples on when grounding emerges in multi-modal large language models, specifically in the Llava 1.5 (7B) variant, using the second probing to prompt the model to segment what is in the referring expression. Figures 12, 14, 13, 15 and 16 show the corresponding predicted masks for the grounding that emerged highlighted in red with the maximum attention point as a black dot. Figure 17 shows the aforementioned five examples with the input prompt, the concept category and the noun phrase corresponding to the best grounding using the oracle selection and the full output text. It clearly shows that the correct output token can correspond to location or color but not necessarily the ground-truth referring expression. While some of the noun phrases and their masks from the point prompting of SAM do correspond to what the noun phrase is describing. It is not always the case, example in Fig. 14 “the flame” was not able to highlight the correct object yet it appeared in the noun phrase corresponding to the location “the top”. It also shows that the majority of correct grounding occurs near the end of the output text after the model reasoned for an answer to the question.
Additionally, this type of mining for attention maps can show potential mistakes that MLLMs fall into while performing the grounding to reason on what is in the image. An example on that is in Fig. 15, where the model potentially mistakes a crack in the wood for the minute hand. Figure 16 shows another clear example, where the MLLM thinks it is seeing a tongue highlighting its potential place near the mouth. Yet the image does not really have the tongue out and consequently the error in the grounding lead to an error in the visual question answering. Thus, our PixFoundation serves as an interesting tool to interpret and understand how MLLMs work and reason to produce the final output with the oracle selection as an upper bound. Yet, there is an open question on how to identify the real mistake from the MLLM vs. simple misalignment between the noun phrase and the grounding emerging, since grounding could emerge with phrases describing it or parts of it instead of the object’s noun phrase itself. We leave this for future work.
the image
a butterfly
orange wings
the dog’s face
the scene
a black and white dog
a black nose
The flame
the match
the top
the image
darkness
The minute hand
the clock
the scene
the 12 o’clock position
The minion
the image
its tongue
a blue overalls
a yellow shirt
| Image | Prompt | Concept Category | Noun Phrase | Output |
| 1 | Identify the butterfly’s wings in the scene. | Color & Appearance | orange wings | In the image, there is a butterfly with orange wings. |
| 3 | Identify the flame of the match in the scene. | Location & Position | the top | The flame of the match is located at the top of the image, surrounded by darkness. |
| 6 | Identify the dog’s face in the scene. | Color & Appearance | a black and white dog | The dog’s face in the scene is a black and white dog with a black nose. |
| 161 | Identify the minute hand of the clock in the scene. | Location & Position | the 12 o’clock position | The minute hand of the clock in the scene is located at the 12 o’clock position. |
| 292 | Identify the minion in the image smiling with its tongue out in the scene. | Color & Appearance | a yellow shirt | The minion in the image is smiling with its tongue out, wearing a blue overalls and a yellow shirt. |
[5pt]
1
[5pt]
3
[5pt]
6
[5pt]
161
[5pt]
292
B.4 PixMMVP benchmark
Figure 18 shows additional results on PixMMVP benchmark comparing different pixel-level MLLMs with our oracle baseline using Llava 1.5 (7B). While GLAMM shows strong pixel-level visual grounding yet we have shown earlier that it is almost incapable of visual question answering which renders the model weak for general purpose tasks. On the other hand, OMG-Llava shows a better balance in pixel-level visual grounding and visual question answering as previously detailed. Nonetheless, the simple mining of attention maps from Llava 1.5 (7B) using the oracle selection which we call PixFoundation shows the strongest capability in both grounding and VQA. In fact, certain MLLMs that were trained with pixel-level visual grounding, such as LISA, have degraded the performance with respect to the hidden information already existing in powerful MLLMs that were not trained with such supervision.
the butterfly’s feet
the window on the school bus
the window on the school bus
the yellow animal’s head
the yellow animal’s head
B.5 PixCV-Bench benchmark
Figure 19 shows qualitative results on PixCV-Bench. It shows that pixel-level MLLMs struggle with segmenting the object annotated by the red box unlike our oracle baseline, PixFoundation. Indeed the attention maps from these MLLMs are looking at the right object annotated by the red box without receiving any pixel-level grounding supervision during training.
cell phone
mouse and keyboard (annotated by the red box)
sports ball and person (annotated by the red box)
chair and cell phone
Appendix C Analysis on the output length
In this section, we provide additional analysis on the output length on average through PixMMVP dataset using the first and second probing schemes. Specifically, we report the output length as the number of characters in the output, and the number of noun phrases extracted from it. The reason to study this, since it has relation to the number of noun phrases and consequently the number of masks our baselines are selecting among. Table 4 shows the average output length computed across PixMMVP dataset, comparing the three base MLLMs. We notice that Cambrian-1 (8B) generates longer outputs with a considerable margin than Llava variants. Hence, we believe the superiority of the oracle upper bound with Cambrian-1 in the grounding has strong correlation to producing longer outputs with more attention maps to mine and select from, than Llava variants. Nonetheless, it makes it more challenging for the automatic baseline which explains its results on PixCV-Bench with respect to the attend and segment. We believe the use of GPT-4o for the automatic selection instead of Cambrian-1 (8B) should improve these results further.
| Model Name | Probing | Output Length | # Noun Phrases |
| Llava 1.5 (7B) | First | 44.2 | 2.3 |
| Llava 1.5 (13B) | First | 45.3 | 2.4 |
| Cambrian-1 (8B) | First | 313.8 | 15.2 |
| Llava 1.5 (7B) | Second | 92.6 | 5.2 |
| Llava 1.5 (13B) | Second | 97.2 | 5.5 |
| Cambrian-1 (8B) | Second | 561.3 | 27.3 |
Appendix D Failure Cases Analysis
In this section, we conduct additional failure case analysis of pixel-level MLLMs and our baselines qualitatively and quantitatively.
D.1 Failures in Visual Question Answering
We start with a fine-grained quantitative analysis of how the studied models perform across PixMMVP and PixCV-Bench. For PixMMVP we follow their scheme to identify the nine visual patterns and report the model’s accuracy with each pattern in Fig. 20. Similarly, we show fine-grained analysis relying on the tasks for the two datasets (ADE20K and COCO) in Fig. 21.
PixMMVP results show that the majority of pixel-level MLLMs, highlighted in blue, suffer in the state, orientation and quantity related tasks. On the other hand, relational context, color and presence of features show the best performance with pixel-level MLLMs. Nonetheless, across all the visual patterns, the MLLMs that were not trained with pixel-level supervision persistently surpass these pixel-level MLLMs with a considerable margin. PixCV-Bench, similarly shows the count task is more challenging than the relational positioning. It also shows that ADE20K dataset serves as a more challenging dataset than COCO.
D.2 Failures in Pixel-level Visual Grounding
Finally, we show qualitatively the failure cases of the oracle upper bound in Fig. 22. It shows failures in segmenting all the object instances in the first row, since the current point prompting assumes one connected component corresponding to each expression. However, certain scenarios, such as the image with the spots on the animal, can lead to these failures in the oracle even when the localisation of some of these is correct. Mechanisms that solve this multi instance scenarios of the same object are left for future work.
Another failure occurring such as in the second row stems from ambiguity in the referring expression itself or failures from SAM identifying the separation between the wall and the ceiling. Hence, the oracle upper bound is generally inheriting SAM failures. However, its main purpose of showing that the hidden information within powerful MLLMs is sufficient to perform pixel-level grounding is achieved, and even surpassing pixel-level MLLMs without degrading their VQA abilities.

### v2

@@
 1 Introduction
-There have been numerous advancements in pixel-level image and video understanding, including tasks such as image/video segmentation (Zhou et al., 2022; Minaee et al., 2021; Kirillov et al., 2023; Ravi et al., 2024), pixel-level visual grounding and reasoning (Rasheed et al., 2024; Lai et al., 2024), depth estimation (Yang et al., 2024) and tracking (Wang et al., 2023).
-The majority of these have been transformed with the emergence of foundation models (Bommasani et al., 2021), specifically multi-modal large language models (MLLMs) (Liu et al., 2023/; Dai et al., 2023).
-Nonetheless, pixel-level MLLMs have shown degradation in their capabilities in chat performance (Lai et al., 2024).
-Recent models tried to address this gap (Zhang et al., 2024b, a), yet they relied on standard evaluation benchmarks overlooking the shortcomings of current MLLMs.
-Recent efforts explored the shortcomings of MLLMs in vision centric benchmarks (Tong et al., 2024b, a).
-Such benchmarks focus on challenging visual patterns and tasks such as counting and relative positioning.
+There have been numerous advancements in pixel-level image and video understanding, including tasks such as image/video segmentation [29, 14, 8, 17], pixel-level visual grounding and reasoning [16, 9], depth estimation [23] and tracking [21].
+The majority of these have been transformed with the emergence of foundation models [1], specifically multi-modal large language models (MLLMs) [11, 3].
+Nonetheless, pixel-level MLLMs have shown degradation in their capabilities in chat performance [9].
+Recent models tried to address this gap [27, 26], yet they relied on standard evaluation benchmarks overlooking the shortcomings of current MLLMs.
+Recent efforts explored the shortcomings of MLLMs in vision centric benchmarks [19, 18].
+Such benchmarks focused on challenging visual patterns and tasks such as counting and relative positioning.
 Nonetheless, these benchmarks did not evaluate the recent pixel-level MLLMs.
-In this work we aim to provide challenging vision centric benchmarks that are dedicated to evaluate these models.
-Through these, we answer the first research question; “Are the current pixel-level MLLMs trained with full grounding supervision heading in the right direction to improve both grounding and visual question answering (VQA) among other capabilities?”.
-Our findings show that the majority of pixel-level MLLMs still fall short in such challenging setting.
-While evidently, some of these show superior performance in visual grounding, we show that MLLMs that were not trained with pixel-level grounding can have better performance without degrading their capabilities.
-There have been recent works showing training-free open vocabulary segmentation emerging from vision language models (Wang et al., 2024; Luo et al., 2024; Hajimiri et al., 2025).
-Concurrent work has specifically explored emerging grounding in MLLMs (Cao et al., 2024).
+In this work we propose challenging vision centric benchmarks that are dedicated to evaluate these models.
+Through these, we answer the first research question; “Are the current pixel-level MLLMs trained with full grounding supervision heading in the right direction to improve both grounding and visual question answering (VQA)?”.
+Our findings show that the majority of pixel-level MLLMs still fall short in such a challenging setting.
+While evidently, some of these show superior performance in visual grounding, we show that MLLMs that were not trained with pixel-level grounding can have better performance without degrading their chat capabilities.
+There have been recent works showing training-free open vocabulary segmentation emerging from vision language models [20, 13, 5].
+Concurrent work has specifically explored emerging grounding in MLLMs [2].
 We are inspired by the previous work to establish a baseline on our proposed benchmarks.
@@
 However, our work documents that emerging grounding in MLLMs does not necessarily coincide with the exact language tokens of the object.
-We show it mostly emerges in the last 40% of the output, and can coincide with concepts about object parts, position/location, color/appearance or context of these objects.
+We show the most frequent emergence in PixMMVP happens in the last 40-60% of the output, and can coincide with concepts about object parts, position, color or context of these objects.
 Fig.
-1 summarizes our two major research questions.
-In summary, our contributions include: (i) Providing pixel-level segmentation annotations and referring expression of the object of interest in the question for two vision centric benchmarks (Tong et al., 2024a, b), which are PixMMVP and PixCV-Bench.
+1 summarizes our research questions.
+In summary, our contributions include: (i) Proposing pixel-level vision centric benchmarks, PixMMVP and PixCV-Bench, with segmentation annotations and referring expression of the object of interest in the corresponding questions [18, 19].
 (ii) Benchmarking recent efforts in pixel-level MLLMs where we show that they degrade VQA capabilities.
@@
 Our mechanism uses the observation that grounding can emerge corresponding to different output tokens describing the object’s appearance or location, not necessarily the exact text of the object of interest.
-We call it PixFoundation, since it mines the foundation model for pixel-level understanding.
+We call it PixFoundation, since it can be plugged with any MLLM to mine for pixel-level understanding.
 2 Related work
 Pixel-level Vision Foundation Models.
-There have been various vision foundation models released that were trained with supervision for the segmentation task (e.g., SAM, SAM 2.0) (Kirillov et al., 2023; Ravi et al., 2024).
-Orthogonal to this, some methods discussed the ability of vision foundation models such as CLIP and BLIP in image segmentation without any segmentation supervision (Luo et al., 2024; Hajimiri et al., 2025; Wang et al., 2024).
+There have been various vision foundation models released that were trained with supervision for the segmentation task (e.g., SAM, SAM 2.0) [8, 17].
+Orthogonal to this, some methods discussed the ability of vision foundation models such as CLIP and BLIP in image segmentation without any segmentation supervision [13, 5, 20].
 Yet, they relied on earlier foundation models that did not incorporate the power of large language models.
-Combining large language models with vision has been extensively researched with pioneering works such as Llava (Liu et al., 2023/, 2024) and instruct-BLIP (Dai et al., 2023).
-Multiple works afterwards focused on pixel-level visual grounding in these MLLMs with full supervision (Lai et al., 2024; Rasheed et al., 2024; Zhang et al., 2024a, b, a, b).
+Combining large language models with vision has been extensively researched with pioneering works such as Llava [11, 12] and instruct-BLIP [3].
+Multiple works afterwards focused on pixel-level visual grounding in these MLLMs with full supervision [9, 16, 26, 27, 26, 27].
 However, these methods were lagging in their chat performance.
-Notably, pixel-level MLLMs were not evaluated on the challenging benchmarks that focused on the shortcomings of MLLMs (Tong et al., 2024b, a).
+Notably, pixel-level MLLMs were not evaluated on the challenging benchmarks that focused on the shortcomings of MLLMs [19, 18].
 Hence, it is still unclear if the pixel-level grounding supervision helped to improve their ability on these challenging tasks or not.
 In this work, we focus on the previous question to have a better understanding of their performance.
-Concurrent work, has shown that without pixel-level supervision there is an emerging ability to perform pixel-level grounding (Cao et al., 2024).
+Concurrent work, has shown that without pixel-level supervision there is an emerging ability to perform pixel-level grounding [2].
 We rely on this method as our baseline, but unlike previous works we provide an insight on when grounding emerges in such MLLMs.
-Benchmarking Multi-modal Large Language Models There is an abundance of standard benchmarks used for evaluating MLLMs (e.g., MMU (Yue et al., 2024)) and their pixel-level counter parts (e.g., refCOCO/+/g (Yu et al., 2016; Kazemzadeh et al., 2014)).
-These benchmarks have pushed the limits on MLLMs capabilities in terms of VQA and visual grounding.
+We propose another baseline that uses a novel and simple mechanism to perform mask selection while taking the previous insight into consideration.
+Our baseline outperforms the previous methods trained with full grounding supervision and the ones without it.
+Benchmarking Multi-modal Large Language Models.
+There is an abundance of standard benchmarks used for evaluating MLLMs (e.g., MMU [25]) and pixel-level benchmarks (e.g., refCOCO/+/g [24, 7]).
+These have pushed the limits on MLLMs capabilities in terms of VQA and visual grounding.
 Nonetheless, there have been various works that discussed the shortcomings of MLLMs.
-In one of these works they discussed specifically CLIP (Radford et al., 2021), which is used in various MLLMs as the visual backbone.
-They proposed a benchmark, MMVP (Tong et al., 2024b), that is focused on the visual aspects within a VQA task designed on CLIP blind pairs.
-More recently, CV-Bench (Tong et al., 2024a) focused on two major tasks that are vision focused which are counting and relative positioning.
+One of them discussed the shortcomings in CLIP [15], which is used in various MLLMs as a visual backbone.
+They proposed a benchmark, MMVP [19], that is focused on the visual aspects within a VQA task.
+More recently, CV-Bench [18] focused on two major tasks that are vision focused which are counting and relative positioning.
 Both were proposed to evaluate MLLMs that do not have the ability to generate segmentation output.
@@
 In this work, we extend these two benchmarks to augment it with pixel-level annotations and referring expressions that correspond to the object of interest within their VQA task.
-These help to understand if failures occurring in these benchmarks stem from grounding or other reasons.
+These help to understand if failures occurring in these benchmarks stem from grounding or VQA capabilities.
 3 Method and benchmarks
@@
 3.1 Benchmarks
-PixMMVP benchmark: We build upon the recently released MMVP (Tong et al., 2024b) which identified clip blind pairs and used them to build a challenging benchmark with the corresponding questions and choices for 300 images.
+PixMMVP benchmark: We build upon the recently released MMVP [19] which identified clip blind pairs and used them to build a challenging benchmark with the corresponding questions and choices for 300 images.
 We augment the aforementioned dataset and manually annotate each question with the corresponding object of interest referring expression, e.g.
@@
 These expressions correspond to what needs to be grounded in the image to answer the question.
-Afterwards, we manually label these objects of interest with polygonal annotations using the VGG annotator (Dutta et al., 2016).
-PixCV-Bench benchmark: For this benchmark we build upon the 2D component of the recently released CV-Bench (Tong et al., 2024a).
-We specifically select the 2D component, since they are sourced from segmentation datasets (i.e., ADE20K (Zhou et al., 2017) and COCO (Lin et al., 2014)), which can be used in our proposed benchmark.
+Afterwards, we manually label these objects of interest with polygonal annotations using the VGG annotator [4].
+PixCV-Bench benchmark: For this benchmark we build upon the 2D component of the recently released CV-Bench [18].
+We specifically select the 2D component, since they are sourced from segmentation datasets (i.e., ADE20K [28] and COCO [10]), which can be used in our proposed benchmark.
 However, the publicly released CV-Bench does not identify the objects in question and their corresponding segmentation.
@@
 Another example that needs manual filtration when the class in question is a broader category than what is inquired upon, e.g., “Pendant Lamp” which is under the category of “Lamp” in ADE20K.
-In such a case we need to filter out the masks of other types such as “Table Lamp”.
-Moreover, we identified missing annotations in rare occasions that required additional intervention.
-Examples include transparent objects such as wine glasses which were partially occluded, yet the ground-truth answer of the CV-Bench question included that object in the count.
-As such, we use VGG annotator for these missing annotations that were not part of the original segmentation masks.
+In such a case, we filter out the masks of other types such as “Table Lamp”.
+Moreover, we identify missing annotations in rare occasions that require additional intervention and manually annotate these missing objects.
 We provide the final PixCV-Bench with referring expressions and their segmentation annotations that can be used to evaluate the grounding ability in relation to the original VQA task.
@@
 First, we highlight the degraded performance in VQA from most of the pixel-level MLLMs that are trained with pixel-level grounding supervision.
-We use for the first prompt “IMGQUESTION?
+We use for that the following prompt, “IMGQUESTION?
 OPTION1 OPTION2…”, as shown in Figure 2.
 Certain pixel-level MLLMs tend to answer the aforementioned question while outputting a corresponding segmentation mask/s for the objects of interest.
-Notably, the worst two models in this task, LISA (Lai et al., 2024) and GLAMM (Rasheed et al., 2024), are not able to provide an answer and rather refer to a segmentation mask.
-On the other hand, OMG-Llava (Zhang et al., 2024b) shows better ability in visual question answering.
+Notably, the worst two models in this task, LISA [9] and GLAMM [16], are not able to provide an answer and rather refer to a segmentation mask.
+On the other hand, OMG-Llava [27] shows better ability in VQA.
 The second shortcoming we discuss is their degraded ability to visually ground objects.
@@
 Figure 3 shows the worst two models in this aspect, which are GLAMM, the region captioning variant, and Llava-G.
-Both fail to segment the specific object in question, while OMG-Llava shows better performance in grounding.
+Both fail to segment the specific object in question, while OMG-Llava shows better performance.
 Third, we highlight another shortcoming, where these MLLMs exhibit degraded ability to follow instructions.
 In order to probe this, we use the following prompt: “IMGQUESTION?
-a.OPTION1 b.OPTION2… Answer with the option’s letter from the given.” Figure 4 shows an example with the answers from the worst two models in this aspect which are LISA (Lai et al., 2024) and Llava-G (Zhang et al., 2024a).
+a.OPTION1 b.OPTION2… Answer with the option’s letter from the given.” Figure 4 shows an example with the answers from the worst two models in this aspect which are LISA [9] and Llava-G [26].
 Both are incapable of following the instruction, yet Llava-G tries to tackle the question unlike LISA.
@@
 In addition to evaluating state-of-the-art pixel-level MLLMs, we propose two baselines and one upper bound.
-The first of which is inspired by a concurrent work (Cao et al., 2024) that identified the emergent grounding in multi-modal large language models without the need for any pixel-level grounding supervision.
+The first of which is inspired by a concurrent work [2] that identified the emergent grounding in multi-modal large language models without the need for any pixel-level grounding supervision.
 Specifically, we use their attend and segment meta architecture as one of our baselines.
@@
 This is further normalized across all the output, for output tokens.
-The attend and segment depends on using spaCy natural language processing tool (Honnibal et al., 2020) to identify the noun phrases and associate them to the ground-truth referring expressions.
+The attend and segment depends on using spaCy natural language processing tool [6] to identify the noun phrases and associate them to the ground-truth referring expressions.
 Thus, the spaCy embeddings closest to the ground-truth expression are used in the mask selection.
-This is followed by extracting the maximum attention point to feed into SAM (Kirillov et al., 2023) as a point prompt.
-| Method | PixGr. Train | MMVP & PixMMVP | ||||
-| Llava 1.5 (7B) (Liu et al., 2023/) | ✗ | 28.7 | 28.0 | - | - | - |
-| Llava 1.5 (13B) (Liu et al., 2023/) | ✗ | 39.3 | 30.0 | - | - | - |
-| Cambrian (8B)* (Tong et al., 2024a) | ✗ | 52.0 | 52.0 | - | - | - |
-| OMG Llava (7B) (Zhang et al., 2024b) | ✓ | 12.0 | 12.0 | 17.8 | 38.0 | 18.2 |
-| GLAMM (7B) (Rasheed et al., 2024) | ✓ | 1.3 | 2.7 | 31.5 | 47.4 | 5.1 |
-| GLAMM - RegCap (7B) (Rasheed et al., 2024) | ✓ | 12.7 | 6.7 | 14.5 | 18.6 | 15.1 |
-| LISA (7B) (Lai et al., 2024) | ✓ | 7.3 | - | 18.1 | 42.9 | 12.5 |
-| Llava-G (7B) (Zhang et al., 2024a) | ✓ | 9.3 | - | 17.8 | 13.5 | 12.2 |
-| Llava 1.5 (7B) + (a+s) (Cao et al., 2024) | ✗ | 28.7 | 28.0 | 11.1 | 11.2 | 16.1 |
-| Llava 1.5 (13B) + (a+s) (Cao et al., 2024) | ✗ | 39.3 | 30.0 | 9.8 | 11.4 | 17.7 |
-| Cambrian (8B)* + (a+s) (Cao et al., 2024) | ✗ | 52.0 | 52.0 | 14.3 | 15.1 | 23.4 |
-| PixFoundation (Llava (7B)) (Ours) | ✗ | 28.7 | 28.0 | 16.9 | 18.8 | 22.7 |
-| PixFoundation (Llava (13B)) (Ours) | ✗ | 39.3 | 30.0 | 14.4 | 18.2 | 24.9 |
-| PixFoundation (Cambrian (8B)*) (Ours) | ✗ | 52.0 | 52.0 | 17.2 | 18.9 | 27.7 |
-| Upper Bound - Oracle Selection | ||||||
-| PixFoundation (Llava (7B)) (Ours) | ✗ | 28.7 | 28.0 | 26.1 | 38.0 | 32.7 |
-| PixFoundation (Llava (13B)) (Ours) | ✗ | 39.3 | 30.0 | 23.6 | 38.2 | 38.7 |
-| PixFoundation (Cambrian (8B)*) (Ours) | ✗ | 52.0 | 52.0 | 52.0 | 56.1 | 54.0 |
+This is followed by extracting the maximum attention point to feed into SAM [8] as a point prompt.
 For our baseline and upper bound, we build upon the previous pipeline and build an oracle upper bound and an automatic baseline.
@@
 IMG”, where EXPR and IMG are the ground-truth expression and the image tokens respectively.
-In our experiments we rely on the open source Cambrian-1(8B) (Tong et al., 2024a) model in our automatic baseline for the mask selection.
-The second modification since SAM has a good understanding of point prompting ambiguity, we process three potential output masks for each prompt instead of a single one.
+In our automatic baseline we rely on GPT-4o for the mask selection.
+The second modification, since SAM has a good understanding of point prompting ambiguity, we process three potential output masks for each prompt instead of one only.
 This enables us to utilize the power of SAM in identifying fine-grained objects and referring expressions that tends to surpass what other MLLMs do, even those trained with pixel-level grounding supervision.
-| Method | PixGr. Train | CV-Bench & PixCV-Bench | ||||
-| Llava 1.5 (7B) (Liu et al., 2023/) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | - | - | - |
-| Llava 1.5 (13B) (Liu et al., 2023/) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | - | - | - |
-| Cambrian (8B)* (Tong et al., 2024a) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | - | - | - |
-| OMG Llava (7B) (Zhang et al., 2024b) | ✓ | 9.2/14.7/12.0 | 36.8/47.4/42.1 | - | 50.5 | 45.9 |
-| GLAMM (7B) (Rasheed et al., 2024) | ✓ | - | - | 30.2 | 51.9 | - |
-| GLAMM - RCap (7B) (Rasheed et al., 2024) | ✓ | 22.8/32.8/27.8 | 46.8/62.0/54.4 | 3.6 | 7.4 | 13.0 |
-| LISA (7B) (Lai et al., 2024) | ✓ | 1.9/5.5/3.7 | - | 16.8 | 48.1 | 6.7 |
-| Llava-G (7B) (Zhang et al., 2024a) | ✓ | 13.9/14.2/14.1 | 5.1/3.7/4.4 | 1.7 | 17.6 | 15.8 |
-| Llava 1.5 (7B) + (a+s) (Cao et al., 2024) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | 5.2 | 15.6 | 24.8 |
-| Llava 1.5 (13B) + (a+s) (Cao et al., 2024) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | 4.7 | 14.9 | 24.0 |
-| Cambrian (8B)* + (a+s) (Cao et al., 2024) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | 18.6 | 15.9 | 29.6 |
-| PixFoundation (Llava (7B)) (Ours) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | 5.3 | 19.1 | 29.0 |
-| PixFoundation (Llava (13B)) (Ours) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | 4.7 | 17.7 | 27.5 |
-| PixFoundation (Cambrian (8B)*)(Ours) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | 12.1 | 16.6 | 27.0 |
-| Upper Bound - Oracle Selection | ||||||
-| PixFoundation (Llava (7B)) (Ours) | ✗ | 16.7/14.7/15.7 | 54.2/66.5/60.4 | 6.3 | 49.7 | 54.5 |
-| PixFoundation (Llava (13B)) (Ours) | ✗ | 14.6/16.7/15.6 | 55.6/67.1/61.3 | 5.3 | 50.6 | 55.4 |
-| PixFoundation (Cambrian (8B)*) (Ours) | ✗ | 55.7/68.7/62.2 | 65.2/79.1/72.2 | 54.3 | 64.4 | 68.1 |
+| Method | PixGr. | PixMMVP | PixCV-Bench | ||||||||
+| Llava 1.5 (7B) [11] | ✗ | 27.3 | 28.0 | - | - | - | 17.4 | 60.3 | - | - | - |
+| Llava 1.5 (13B) [11] | ✗ | 39.3 | 30 | - | - | - | 14.5 | 61.4 | - | - | - |
+| Cambrian (8B)* [18] | ✗ | 52.0 | 52.0 | - | - | - | 62.2 | 72.2 | - | - | - |
+| OMG Llava (7B)** [27] | ✓ | 12.0 | 12.0 | 17.8 | 38.0 | 18.2 | 12.0 | 42.1 | - | 50.5 | 45.9 |
+| GLAMM (7B) [16] | ✓ | 1.3 | 2.7 | 31.5 | 47.4 | 5.1 | - | - | 30.2 | 51.9 | - |
+| GLAMM - RegCap (7B) [16] | ✓ | 12.7 | 6.7 | 14.5 | 18.6 | 15.1 | 27.8 | 54.4 | 3.6 | 7.4 | 13.0 |
+| LISA (7B) [9] | ✓ | 7.3 | - | 18.1 | 42.9 | 12.5 | 3.7 | - | 16.8 | 48.1 | 6.7 |
+| Llava-G (7B) [26] | ✓ | 9.3 | - | 17.8 | 13.5 | 12.2 | 14.1 | 4.4 | 1.7 | 17.6 | 15.8 |
+| Llava 1.5 (7B) + (a+s) [2] | ✗ | 27.3 | 28.0 | 11.1 | 11.2 | 16.0 | 17.4 | 60.3 | 5.2 | 15.7 | 24.9 |
+| Llava 1.5 (13B) + (a+s) [2] | ✗ | 39.3 | 30 | 9.8 | 11.4 | 17.7 | 14.5 | 61.4 | 4.7 | 14.9 | 24.0 |
+| Cambrian (8B)* + (a+s) [2] | ✗ | 52.0 | 52.0 | 14.3 | 15.1 | 23.4 | 62.2 | 72.2 | 18.6 | 15.9 | 29.6 |
+| PixFoundation (7B) (Ours) | ✗ | 27.3 | 28.0 | 18.8 | 25.9 | 26.9 | 17.4 | 60.3 | 5.4 | 28.5 | 38.7 |
+| PixFoundation (13B) (Ours) | ✗ | 39.3 | 30 | 16.9 | 25.0 | 30.6 | 14.5 | 61.4 | 4.8 | 27.6 | 38.1 |
+| PixFoundation (8B)* (Ours) | ✗ | 52.0 | 52.0 | 29.6 | 30.3 | 38.3 | 62.2 | 72.2 | 23.9 | 33.1 | 45.4 |
+| Upper Bound - Oracle Selection | |||||||||||
+| PixFoundation (7B) (Ours) | ✗ | 27.3 | 28.0 | 26.1 | 38.0 | 32.2 | 17.4 | 60.3 | 6.3 | 49.7 | 54.5 |
+| PixFoundation (13B) (Ours) | ✗ | 39.3 | 30 | 23.6 | 38.2 | 38.7 | 14.5 | 61.4 | 5.3 | 50.6 | 55.5 |
+| PixFoundation (8B)* (Ours) | ✗ | 52.0 | 52.0 | 52.0 | 56.1 | 54.0 | 62.2 | 72.2 | 54.3 | 64.4 | 68.1 |
 the dorsal fin of the animal
+| Image | Prompt | Concept Category | Noun Phrase | Output |
+| 1 | Identify the butterfly’s wings in the scene. | Color & Appearance | orange wings | In the image, there is a butterfly with orange wings. |
+| 3 | Identify the flame of the match in the scene. | Location & Position | the top | The flame of the match is located at the top of the image, surrounded by darkness. |
+| 6 | Identify the dog’s face in the scene. | Color & Appearance | a black and white dog | The dog’s face in the scene is a black and white dog with a black nose. |
+| 161 | Identify the minute hand of the clock in the scene. | Location & Position | the 12 o’clock position | The minute hand of the clock in the scene is located at the 12 o’clock position. |
+[5pt]
+1
+[5pt]
+3
+[5pt]
+6
+[5pt]
+161
 4 Experiments
@@
 On each benchmark we evaluate the VQA and visual grounding capabilities following three probing techniques and reporting their metrics.
-The first probing is to evaluate the VQA ability, where the accuracy is computed using GPT-4o following (Tong et al., 2024b) as, .
+The first probing is to evaluate the VQA ability, where the accuracy is computed using GPT-4o following [19] as, .
 If the model generates a segmentation without explicitly asking it to, it is evaluated with respect to the ground-truth referring segmentation in terms of mean intersection over union as, .
 The second probing prompts the model to identify the referred expression then evaluates the mean intersection over union reported as, .
-The third probing following (Tong et al., 2024a) instructs the model to generate a single option letter and evaluate the accuracy directly without GPT-4o, reported as, .
+The third probing following [18] instructs the model to generate a single option letter and evaluate the accuracy directly without GPT-4o, reported as, .
 There is a need for the first probing since some of the recent pixel-level MLLMs face challenges in following instructions.
 We evaluate the score of each model, , which is the harmonic mean across the maximum of both pixel-level visual grounding and VQA,
-We mainly focus on evaluating four state-of-the-art pixel-level MLLMs; LISA (Lai et al., 2024), GLAMM (Rasheed et al., 2024), OMG-Llava (Zhang et al., 2024b) and Llava-G (Zhang et al., 2024a).
+We mainly focus on evaluating four state-of-the-art pixel-level MLLMs; LISA [9], GLAMM [16], OMG-Llava [27] and Llava-G [26].
 For GLAMM we use two variants; the original model (GLAMM) and the one fine-tuned for region-level captioning, (GLAMM-RegCap).
-We use Hugging Face for these models weights, refer to Appendix A.
+For details on the models weights, refer to App.
+A.
 Baselines and upper bound implementation details.
 We evaluate: (i) the attend and segment (a+s), (ii) the oracle selection relying on the highest intersection over union in selecting the correctly predicted masks (PixFoundation), and (iii) the automatic selection, (PixFoundation).
-These are implemented on top of three base MLLMs which are, Llava 1.5 (7B, 13B) (Liu et al., 2024) and Cambrian-1(8B) (Tong et al., 2024a).
-The automatic selection is implemented using the Cambrian-1 (8B) model, yet it can be further improved using GPT-4o, refer to Appendix A for more details.
+These are implemented on top of three base MLLMs which are, Llava 1.5 (7B, 13B) [12] and Cambrian-1(8B) [18].
+The automatic selection is implemented using GPT-4o, refer to App.
+A for more details.
 4.2 Are the current pixel-level MLLMs heading in the right direction?
-In order to answer this, we evaluate each of these pixel-level MLLMs capability in VQA in challenging tasks.
+In order to answer this, we evaluate each of these pixel-level MLLMs VQA capabilities in challenging tasks.
 Additionally, we evaluate their ability to visually ground the objects of interest in these questions.
-PixMMVP Benchmark.
-Table 1 shows the results on the challenging PixMMVP benchmark.
+Table 1 shows the results on the challenging PixMMVP and PixCV-Bench.
 From the accuracy of VQA, MLLMs that are not trained with pixel-level grounding surpass their pixel-level counterpart with up to 16%.
-The best in pixel-level MLLMs score in this aspect is GLAMM-RegCap (Zhang et al., 2024b) yet it has degraded ability to generate segmentation.
-On the other hand, when looking at pixel-level visual grounding we find the best model, GLAMM (Rasheed et al., 2024), has a weak ability in VQA or following instructions.
-Moreover, it shows LISA and Llava-G are incapable of following the instruction to output the option letter reported in, .
-Looking at the bottom three rows, the oracle confirms that MLLMs that were never trained with pixel-level grounding have the correct grounding within their learned attention maps, refer to Fig.
-5.
-Additional qualitative analysis is in Appendix B.
-Looking at the final score, , the oracle variant, PixFoundation (7B), outperforms the corresponding best pixel-level MLLMs, OMG-Llava (7B), by a considerable margin while the automatic outperforms it with up to 4%.
-While the attend and segment baseline (Cao et al., 2024) lags behind our automatic method with more than 6%.
-PixCV-Bench Benchmark.
-Table 2 shows the results on PixCV-Bench.
-Persistently all the pixel-level MLLMs still lag behind the MLLMs that were never trained with pixel grounding supervision in VQA.
-Nonetheless, OMG-Llava strikes the right balance in both VQA and pixel-level grounding with the highest score, , within the 7B models.
-While, GLAMM, shows the best pixel-level grounding yet it is incapable of VQA on this benchmark.
-Looking at the bottom three rows for the oracle similar findings to PixMMVP emerge.
-The attend and segment baseline and the automatic PixFoundation baseline, both lag behind the oracle as expected yet they show competitive performance in pixel-level visual grounding, even better than some of the MLLMs trained with pixel-level supervision.
-Our automatic PixFoundation outperforms the attend and segment using the Llava variants, and is competitive in Cambrian-1.
-Cambrian-1 is more challenging as it generates more tokens, thus selecting among higher number of masks, refer to Appendix C.
+The best in pixel-level MLLMs score in this aspect is GLAMM-RegCap [27] yet it has degraded ability to generate segmentation.
+On the other hand, when looking at pixel-level visual grounding we find the best model, GLAMM [16], has a weak ability in VQA or following instructions.
+Moreover, it shows LISA and Llava-G are mostly incapable of following the instruction to output the option letter reported in, .
+OMG-Llava strikes the right balance in both VQA and pixel-level grounding with the highest score, , within pixel-level MLLMs.
+However, looking at the bottom three rows, the oracle, PixFoundation, confirms that MLLMs that were never trained with pixel-level grounding have the correct grounding within their learned attention maps yielding higher score, , than OMG-Llava.
+Qualitative comparison is provided in Fig.
+5 and additional examples are in App.
+B.
+Additionally, the automatic baseline outperforms OMG-Llava in PixMMVP with a considerable margin and is on-par to it on PixCV-Bench.
+Furthermore, the attend and segment baseline [2] lags behind our automatic method with more than 10% in both datasets.
+Refer to App.
+C for a failure case analysis.
+Finally, we evaluate whether the failures of these MLLMs occur in visual grounding, VQA or both.
+Figure 6 shows the frequency of failures per category, where the majority stem from failures in both especially in the pixel-level MLLMs.
+While the vanilla MLLMs perform better in the VQA and tend to fail more in the grounding task.
 Summary.
 In summary, pixel-level grounding supervision degrades MLLMs ability in VQA and sometimes even their generalization in grounding.
-Failure case analysis is provided in Appendix D.
 We show that MLLMs trained with pixel-level supervision lag behind vanilla MLLMs using simple mechanisms to extract grounding, and the oracle indicates there is an opportunity to improve this.
-Moreover, we show that grounding might not coincide with the noun phrase most similar to the referred expression, where our oracle upper bound is surpassing the attend and segment.
+Moreover, we show that grounding might not coincide with the noun phrase most similar to the referred expression, where our oracle upper bound and automatic baseline are both surpassing the attend and segment.
 4.3 When does grounding emerge in MLLMs?
@@
 Accordingly, Fig.
-6(a) shows the location percentages histogram, binned at 10%, for the three base MLLMs reporting the oracle selection and evaluating on PixMMVP benchmark using the second probing.
-In the Llava 1.5 variants, grounding mostly emerges at the last 40% of the output text, while for Cambrian it is mostly at the last 10%.
+7(a) shows the location percentages histogram, binned at 10%, for the three base MLLMs reporting the oracle selection and evaluating on PixMMVP benchmark using the second probing.
+In the Llava 1.5 variants, the highest grounding emerges at the last 40% of the output text, while for Cambrian it emerges at the last 60%.
 When - concept.
 For the second analysis we look into the concept category that the correct output word/phrase corresponds to.
-The previous assumption in other works is that grounding emerges in exact noun or noun phrase of the object of interest.
+The previous assumption in other works is that grounding emerges in the exact noun/noun phrase of the object of interest.
 Except our analysis confirms that this is not necessarily the case.
 We take the correct noun/noun phrase where the grounding emerges based on the oracle from all the three variants, then we pass it to GPT-4o to request a grouping of these concepts.
-Specifically, we prompt it with the following “Can you group these noun phrases into categories of abstract concepts.
-You are provided with pairs of (noun phrase, its full context).”, where the entire output text is the full context.
-After manual filtration and merging it resulted into six main groups, which are: (i) Color and appearance, (ii) location and position, (iii) object parts, (iv) context and setting, (v) objects and entities, and (vi) State.
+It resulted into six main groups, which are: (i) Color and appearance, (ii) location and position, (iii) object parts, (iv) context and setting, (v) objects and entities, and (vi) State.
 We then prompt for each of the noun/noun phrase, GPT-4o, to categorize it within these six categories.
 The histogram of the occurrences of these concept categories is shown in Fig.
-6(b).
-It clearly conveys that in certain scenarios the correct output when grounding emerges can be describing the position or the location of the object of interest not necessarily the exact ground-truth referring expression.
-While the majority occurs under “Objects and Entities”, there are still percentage of output coinciding with position, color, context, object part, or even the state.
-Further qualitative results are provided in the Appendix B.
+7(b).
+It clearly conveys that in certain scenarios the correct output when grounding emerges can be describing the position or the color of the object of interest not necessarily the exact ground-truth referring expression.
+Fig.
+8 shows qualitative examples of these scenarios from PixMMVP using the second probing.
+Additional results are provided in App.
+B.
 Random vs.
@@
 Summary.
-In summary, we found that grounding emerges in the last 40% of the output text which might indicate that the reasoning of the model to provide a response impacts when the correct visual grounding can occur.
+In summary, we found that highest frequent emergence of grounding is in the last 40-60% of the output text on PixMMVP which might indicate that the reasoning of the model on this challenging task to provide a response impacts when the correct visual grounding can occur.
 More importantly, we show that grounding in MLLMs can emerge in the noun phrase that corresponds to color, position or other characteristics of the object of interest and not necessarily the exact referring expression.
@@
 Models.
-We also detail the model checkpoints we use for the four pixel-level MLLMs and their variants, retrieved from HuggingFace (Wolf et al., 2019) in Table 3.
+We also detail the model checkpoints we use for the four pixel-level MLLMs and their variants, retrieved from HuggingFace [22] in Table 2.
 These also include the model checkpoints used for the base MLLMs that were not trained with pixel-level visual grounding.
 It is worth noting, that for GLAMM we use two variants (FullScope and RegCap) since their base model (i.e., FullScope) has low performance in the visual question answering task.
-As such we use the other variant for GLAMM that was fine-tuned for region-level captioning using RefCOCOg and Visual Genome dataset.
-Furthermore, we provide details on the oracle selection mechanism, where in the case of multiple masks exhibiting the same IoU we select the last occurring one in the when analysis, since we need to select the location of when it occurs.
-While in the quantitative and qualitative evaluation, we resort to simply not selecting any mask.
-These occur in few cases when the ground-truth mask is all background, thus any mask IoU will result in zero.
+As such we use the other variant for GLAMM that was fine-tuned for region-level captioning using RefCOCOg dataset.
+Furthermore, we provide details on the oracle selection mechanism, we discard the cases where the ground-truth segmentation is all background in the when analysis, since there is no ground-truth grounding to evaluate against.
+These occur in few cases in PixMMVP.
 Additionally, we provide details on the SAM model that is used in the three baselines and upper bounds in our benchmarks, where we use the ViT-H variant.
 Finally, we provide an illustrative example of our automatic selection mechanism with the corresponding predictions on PixMMVP using Llava 1.5 (7B) in Fig.
-9.
-Our automatic selection goes through an iterative process of prompting the selected MLLM, in our case Cambrian-1, with three images overlayed with predicted segmentation to select the best within each group of three.
+11.
+Our automatic selection goes through an iterative process of prompting the selected MLLM, in our case GPT-4o, with groups of images highlighting the predicted segmentation to select the best within each group.
 In the final stage the best images are used to prompt the MLLM to select the final selected mask that best describes the object of interest.
-In the oracle upper bound whenever the model is to be evaluated in a multiple object scenario we take all the possible pairs of the masks then select the best pair based on the highest intersection over union.
 | Model Name | Model Checkpoint |
@@
 | Llava-G | Haozhangcx/llava_grounding_gd_vp |
-| Llava 1.5 (7B) | 4bit/llava-v1.5-7b |
-| Llava 1.5 (13B) | 4bit/llava-v1.5-13b-3GB |
+| Llava 1.5 (7B) | liuhaotian/llava-v1.5-7b |
+| Llava 1.5 (13B) | liuhaotian/llava-v1.5-13b |
 | Cambrian-1 (8B) | nyu-visionx/cambrian-8b |
 Evaluation.
-We also provide the details on computing the visual question answering accuracy using GPT-4o in the first protocol (Tong et al., 2024b).
+We also provide the details on computing the visual question answering accuracy using GPT-4o in the first protocol [19].
 We use the following prompt: “Given the following question QUESTION, the correct answer is ANSWER.
@@
 Additionally, it shows that our simple automatic mechanism, PixFoundation, surpasses the attend and segment as well on PixMMVP.
-Attend and Segment (Cao et al., 2024) (Cambrian-1 (8B))
+Attend and Segment [2] (Cambrian-1 (8B))
 PixFoundation (Cambrian-1 (8B))
@@
 In this section, we visualize the normalized attention maps, , in Fig.
-11.
+13.
 We show two examples for Cambrian-1 (8B) from PixMMVP using the first probing where we directly prompt the model with question and options.
@@
 We show additional examples on when grounding emerges in multi-modal large language models, specifically in the Llava 1.5 (7B) variant, using the second probing to prompt the model to segment what is in the referring expression.
-Figures 12, 14, 13, 15 and 16 show the corresponding predicted masks for the grounding that emerged highlighted in red with the maximum attention point as a black dot.
-Figure 17 shows the aforementioned five examples with the input prompt, the concept category and the noun phrase corresponding to the best grounding using the oracle selection and the full output text.
+Figures 14, 16, 15 and 17 show the corresponding predicted masks for the grounding that emerged highlighted in red with the maximum attention point as a black dot.
+Figure 8 in the main text shows the aforementioned four examples with the input prompt, the concept category and the noun phrase corresponding to the best grounding using the oracle selection and the full output text.
 It clearly shows that the correct output token can correspond to location or color but not necessarily the ground-truth referring expression.
@@
 An example on that is in Fig.
-15, where the model potentially mistakes a crack in the wood for the minute hand.
-Figure 16 shows another clear example, where the MLLM thinks it is seeing a tongue highlighting its potential place near the mouth.
-Yet the image does not really have the tongue out and consequently the error in the grounding lead to an error in the visual question answering.
+17, where the model potentially mistakes a crack in the wood for the minute hand.
 Thus, our PixFoundation serves as an interesting tool to interpret and understand how MLLMs work and reason to produce the final output with the oracle selection as an upper bound.
@@
 the 12 o’clock position
-The minion
-the image
-its tongue
-a blue overalls
-a yellow shirt
-| Image | Prompt | Concept Category | Noun Phrase | Output |
-| 1 | Identify the butterfly’s wings in the scene. | Color & Appearance | orange wings | In the image, there is a butterfly with orange wings. |
-| 3 | Identify the flame of the match in the scene. | Location & Position | the top | The flame of the match is located at the top of the image, surrounded by darkness. |
-| 6 | Identify the dog’s face in the scene. | Color & Appearance | a black and white dog | The dog’s face in the scene is a black and white dog with a black nose. |
-| 161 | Identify the minute hand of the clock in the scene. | Location & Position | the 12 o’clock position | The minute hand of the clock in the scene is located at the 12 o’clock position. |
-| 292 | Identify the minion in the image smiling with its tongue out in the scene.
-| Color & Appearance | a yellow shirt | The minion in the image is smiling with its tongue out, wearing a blue overalls and a yellow shirt.
-|
-[5pt]
-1
-[5pt]
-3
-[5pt]
-6
-[5pt]
-161
-[5pt]
-292
 B.4 PixMMVP benchmark
@@
 chair and cell phone
-Appendix C Analysis on the output length
-In this section, we provide additional analysis on the output length on average through PixMMVP dataset using the first and second probing schemes.
-Specifically, we report the output length as the number of characters in the output, and the number of noun phrases extracted from it.
-The reason to study this, since it has relation to the number of noun phrases and consequently the number of masks our baselines are selecting among.
-Table 4 shows the average output length computed across PixMMVP dataset, comparing the three base MLLMs.
-We notice that Cambrian-1 (8B) generates longer outputs with a considerable margin than Llava variants.
-Hence, we believe the superiority of the oracle upper bound with Cambrian-1 in the grounding has strong correlation to producing longer outputs with more attention maps to mine and select from, than Llava variants.
-Nonetheless, it makes it more challenging for the automatic baseline which explains its results on PixCV-Bench with respect to the attend and segment.
-We believe the use of GPT-4o for the automatic selection instead of Cambrian-1 (8B) should improve these results further.
-| Model Name | Probing | Output Length | # Noun Phrases |
-| Llava 1.5 (7B) | First | 44.2 | 2.3 |
-| Llava 1.5 (13B) | First | 45.3 | 2.4 |
-| Cambrian-1 (8B) | First | 313.8 | 15.2 |
-| Llava 1.5 (7B) | Second | 92.6 | 5.2 |
-| Llava 1.5 (13B) | Second | 97.2 | 5.5 |
-| Cambrian-1 (8B) | Second | 561.3 | 27.3 |
-Appendix D Failure Cases Analysis
+Appendix C Failure case analysis
 In this section, we conduct additional failure case analysis of pixel-level MLLMs and our baselines qualitatively and quantitatively.
-D.1 Failures in Visual Question Answering
+C.1 Failures in visual question answering
 We start with a fine-grained quantitative analysis of how the studied models perform across PixMMVP and PixCV-Bench.
@@
 It also shows that ADE20K dataset serves as a more challenging dataset than COCO.
-D.2 Failures in Pixel-level Visual Grounding
+C.2 Failures in pixel-level visual grounding
 Finally, we show qualitatively the failure cases of the oracle upper bound in Fig.
@@
 However, its main purpose of showing that the hidden information within powerful MLLMs is sufficient to perform pixel-level grounding is achieved, and even surpassing pixel-level MLLMs without degrading their VQA abilities.
+Appendix D Additional quantitative analysis
+D.1 When grounding emerges - CV-Bench
+In Fig.
+23(a) we show the analysis on when grounding emerges on PixCV-Bench in terms of the frequency of the grounding location.
+It is worth noting that PixMMVP is more challenging than PixCV-Bench evidently from the reported IoU and accuracy metrics on both reported in Table 1.
+It seems on the less challenging dataset PixCV-Bench, grounding tends to emerge frequently near the beginning of the output.
+This might relate to PixMMVP being more challenging in terms of the level of reasoning than CV-Bench or the fact that PixMMVP poses a harder referring segmentation task than PixCV-Bench, which is mostly using the class names.
+However, the consistent finding among both datasets is that grounding can emerge coinciding to various concept categories whether location, color or state as shown in Fig.
+23(b).
+D.2 Analysis on the output length
+In this section, we provide additional analysis on the output length on average through PixMMVP dataset using the first and second probing schemes.
+Specifically, we report the output length as the number of characters in the output, and the number of noun phrases extracted from it.
+The reason to study this, since it has relation to the number of noun phrases and consequently the number of masks our baselines are selecting among.
+Table 3 shows the average output length computed across PixMMVP dataset, comparing the three base MLLMs.
+We notice that Cambrian-1 (8B) generates longer outputs with a considerable margin than Llava variants.
+Hence, we believe the superiority of the oracle upper bound with Cambrian-1 in the grounding has strong correlation to producing longer outputs with more attention maps to mine and select from, than Llava variants.
+| Model Name | Probing | Output Length | # Noun Phrases |
+| Llava 1.5 (7B) | First | 44.2 | 2.3 |
+| Llava 1.5 (13B) | First | 45.3 | 2.4 |
+| Cambrian-1 (8B) | First | 313.8 | 15.2 |
+| Llava 1.5 (7B) | Second | 92.6 | 5.2 |
+| Llava 1.5 (13B) | Second | 97.2 | 5.5 |
+| Cambrian-1 (8B) | Second | 561.3 | 27.3 |

### v4

@@
 Abstract
-Multiple works have emerged to push the boundaries on multi-modal large language models (MLLMs) towards pixel-level understanding.
-Such approaches have shown strong performance on benchmarks for referring expression segmentation and grounded conversation generation.
-The current trend in pixel-level MLLMs is to train with pixel-level grounding supervision on large-scale labelled data.
-However, we show that such MLLMs when evaluated on recent challenging vision centric benchmarks, exhibit a weak ability in visual question answering.
-Surprisingly, some of these methods even downgrade the grounding ability of MLLMs that were never trained with such supervision.
-In this work, we propose two novel challenging benchmarks and show that MLLMs without pixel-level grounding supervision can outperform the state of the art in such tasks when evaluating both the pixel-level grounding and visual question answering.
-We propose simple baselines to extract the grounding information that can be plugged into any MLLM, which we call as PixFoundation.
-More importantly, we study the research question of “When does grounding emerge in MLLMs that are not trained with pixel-level grounding supervision?” We show that grounding can coincide with object parts or location/appearance information.
-Code repository is at https://github.com/MSiam/PixFoundation/.
+Multiple works have emerged to push the boundaries of multi-modal large language models (MLLMs) towards pixel-level understanding.
+The current trend is to train MLLMs with pixel-level grounding supervision in terms of masks on large-scale labelled data and specialized decoders for the segmentation task.
+However, we show that such MLLMs when evaluated on recent challenging vision-centric benchmarks, exhibit a weak ability in visual question answering (VQA).
+Surprisingly, some of these methods even downgrade the grounding ability of MLLMs that were never trained with such pixel-level supervision.
+In this work, we propose two novel challenging benchmarks with paired evaluation for both VQA and grounding.
+We demonstrate that simple baselines that are not unified achieve performance that matches or surpasses some of the pixel-level MLLMs.
+Our paired benchmarks and evaluation enable additional analysis on the reasons for failure with respect to VQA and/or grounding.
+Furthermore, we propose a prompt sensitivity analysis on both the language and visual prompts tailored for the grounding task.
+More importantly, we study the research question of “When does grounding emerge in MLLMs with respect to the output tokens?” We propose an interpretability tool that can be plugged into any MLLM to study the aforementioned question.
+We show that grounding does not necessarily coincide with the exact referring expression in the output, but can coincide with the object parts, its location, appearance, context or state.
+Code and datasets are publicly available at https://msiam.github.io/PixFoundationSeries/.
 1 Introduction
-There have been numerous advancements in pixel-level image and video understanding, including tasks such as image/video segmentation [29, 14, 8, 17], pixel-level visual grounding and reasoning [16, 9], depth estimation [23] and tracking [21].
-The majority of these have been transformed with the emergence of foundation models [1], specifically multi-modal large language models (MLLMs) [11, 3].
-Nonetheless, pixel-level MLLMs have shown degradation in their capabilities in chat performance [9].
-Recent models tried to address this gap [27, 26], yet they relied on standard evaluation benchmarks overlooking the shortcomings of current MLLMs.
-Recent efforts explored the shortcomings of MLLMs in vision centric benchmarks [19, 18].
-Such benchmarks focused on challenging visual patterns and tasks such as counting and relative positioning.
-Nonetheless, these benchmarks did not evaluate the recent pixel-level MLLMs.
-In this work we propose challenging vision centric benchmarks that are dedicated to evaluate these models.
-Through these, we answer the first research question; “Are the current pixel-level MLLMs trained with full grounding supervision heading in the right direction to improve both grounding and visual question answering (VQA)?”.
+There have been numerous advancements in pixel-level understanding, including semantic, instance, panoptic segmentation and their video counterpart [36, 18, 11, 21].
+In addition, significant progress has been made in visual grounding and reasoning [20, 12], depth estimation [30] and pixel-level tracking [27].
+Visual grounding is tied to relating the visual content with language; examples of its pixel-level tasks include referring segmentation and grounded conversation generation.
+The majority of these models and tasks have been transformed with the emergence of foundation models [2], especially multi-modal large language models (MLLMs) [15, 5].
+Nonetheless, pixel-level MLLMs, i.e., MLLMs designed for pixel-level tasks, have shown degradation in their language capabilities [12].
+Recent efforts explored the shortcomings of standard MLLMs in vision-centric benchmarks [24, 23].
+Such benchmarks focused on challenging visual tasks that require a form of grounding, such as counting.
+Nonetheless, these benchmarks did not evaluate the recent pixel-level ones.
+In this work, we focus on pixel-level visual grounding with a segmentation mask output and propose challenging vision-centric benchmarks that are dedicated to evaluating pixel-level MLLMs, which we call PixMMVP and PixCV-Bench.
+We then provide a comprehensive paired evaluation for VQA and grounding.
+Our paired evaluation means that the referring segmentation is related to the object of interest in the visual question.
+Through these, we answer the first research question; “Are the current pixel-level MLLMs trained with mask supervision heading in the right direction to improve both grounding and visual question answering (VQA)?”.
 Our findings show that the majority of pixel-level MLLMs still fall short in such a challenging setting.
-While evidently, some of these show superior performance in visual grounding, we show that MLLMs that were not trained with pixel-level grounding can have better performance without degrading their chat capabilities.
-There have been recent works showing training-free open vocabulary segmentation emerging from vision language models [20, 13, 5].
-Concurrent work has specifically explored emerging grounding in MLLMs [2].
-We are inspired by the previous work to establish a baseline on our proposed benchmarks.
-However, unlike their method we focus on the second research question of “When does grounding emerge in MLLMs that are not trained with full grounding supervision?”.
-Previous efforts studying emergent segmentation and/or grounding assumed correspondence with the specific language tokens (i.e., nouns or noun phrases) of the objects of interest.
-However, our work documents that emerging grounding in MLLMs does not necessarily coincide with the exact language tokens of the object.
-We show the most frequent emergence in PixMMVP happens in the last 40-60% of the output, and can coincide with concepts about object parts, position, color or context of these objects.
-Fig.
-1 summarizes our research questions.
-In summary, our contributions include: (i) Proposing pixel-level vision centric benchmarks, PixMMVP and PixCV-Bench, with segmentation annotations and referring expression of the object of interest in the corresponding questions [18, 19].
-(ii) Benchmarking recent efforts in pixel-level MLLMs where we show that they degrade VQA capabilities.
-More importantly, some of them lag in visual grounding with respect to simple techniques of extracting the segmentation from vanilla MLLMs, i.e., MLLMs that are not trained for pixel-level grounding.
-(iii) We provide a simple mechanism for extracting segmentation from vanilla MLLMs, with an understanding of when grounding emerges.
+While evidently, some of these show superior performance in grounding, we show that such MLLMs with specialized segmentation decoders in their architecture and with mask supervision can be worse than their counterpart baselines when considering both visual and language abilities.
+There has been concurrent work [29] that observed the degradation of pixel-level MLLMs’ VQA abilities.
+Nonetheless, previous efforts used standard evaluation benchmarks that evaluate each task separately.
+Our benchmarks provide a paired evaluation designed to be vision-centric, with a focus on what MLLMs fall short in and provide the means to interpret the failures of these MLLMs and whether they are stemming from grounding, language capabilities or both.
+More importantly, unlike concurrent efforts, we focus on the second question of “When does grounding emerge in MLLMs with respect to the output tokens?” Our work documents that emerging grounding in MLLMs does not necessarily coincide with the exact language tokens of the referred object, as shown in Fig.
+1.
+We design an interpretability tool that uses the maximum attention point with respect to the output tokens and analyze the ones that correspond to the best emergent point grounding, using quantitative and qualitative means.
+While recent works show training-free segmentation emerging from vision language models corresponding to the referred expression [25, 16, 7, 3], our study quantifies that, within MLLMs, that is not necessarily the case.
+In summary, our contributions include: (i) Proposing paired pixel-level vision-centric benchmarks, PixMMVP and PixCV-Bench, with segmentation annotations and referring expression of the object of interest in the corresponding questions.
+In addition to proposing prompt sensitivity variations tailored to the grounding task.
+(ii) Benchmarking recent efforts in pixel-level MLLMs where we show that the majority degrade VQA capabilities.
+More importantly, some of them lag in visual grounding with respect to baselines that are not unified.
+(iii) We propose a novel interpretability mechanism to study when grounding emerges in MLLMs with respect to the output tokens that benefit from our paired benchmarks.
 Our mechanism uses the observation that grounding can emerge corresponding to different output tokens describing the object’s appearance or location, not necessarily the exact text of the object of interest.
-We call it PixFoundation, since it can be plugged with any MLLM to mine for pixel-level understanding.
 2 Related work
 Pixel-level vision foundation models.
-There have been various vision foundation models released that were trained with supervision for the segmentation task (e.g., SAM, SAM 2.0) [8, 17].
+There have been various vision foundation models trained for the segmentation task (e.g., SAM and SAM 2.0) [11, 21].
 Orthogonal to this, some methods discussed the ability of vision foundation models such as CLIP and BLIP in image segmentation without any segmentation supervision [16, 7, 25].
@@
 Combining large language models with vision has been extensively researched with pioneering works such as LLaVA [15, 14] and instruct-BLIP [5].
-Multiple works afterwards focused on pixel-level visual grounding in these MLLMs with full supervision [9, 16, 26, 27, 26, 27].
+Multiple works afterwards focused on pixel-level visual grounding in these MLLMs with mask supervision and specialized segmentation decoders [12, 20, 33, 34].
 However, these methods were lagging in their chat performance.
 Notably, pixel-level MLLMs were not evaluated on the challenging benchmarks that focused on the shortcomings of MLLMs [24, 23].
-Hence, it is still unclear if the pixel-level grounding supervision helped to improve their ability on these challenging tasks or not.
+Hence, it is still unclear if mask supervision helped to improve their grounding ability on these challenging tasks.
 In this work, we focus on the previous question to have a better understanding of their performance.
-Concurrent work, has shown that without pixel-level supervision there is an emerging ability to perform pixel-level grounding [2].
-We rely on this method as our baseline, but unlike previous works we provide an insight on when grounding emerges in such MLLMs.
-We propose another baseline that uses a novel and simple mechanism to perform mask selection while taking the previous insight into consideration.
-Our baseline outperforms the previous methods trained with full grounding supervision and the ones without it.
+While concurrent work [26] designed better pixel-level MLLMs that are capable of strong VQA and grounding, we study it in relation to its baseline standard MLLM that was not trained with mask supervision and compare their robustness within a language and visual prompt sensitivity evaluation.
 Benchmarking multi-modal large language models.
-There is an abundance of standard benchmarks used for evaluating MLLMs (e.g., MMU [25]) and pixel-level benchmarks (e.g., refCOCO/+/g [24, 7]).
+There is an abundance of standard benchmarks used for evaluating MLLMs (e.g., MMU [32]) and visual grounding benchmarks (e.g., refCOCO/+/g [31, 10]).
 These have pushed the limits on MLLMs capabilities in terms of VQA and visual grounding.
@@
 They proposed a benchmark, MMVP [24], that is focused on the visual aspects within a VQA task.
-More recently, CV-Bench [18] focused on two major tasks that are vision focused which are counting and relative positioning.
+More recently, CV-Bench [23] focused on two major tasks that are vision-focused, which are counting and relative positioning.
 Both were proposed to evaluate MLLMs that do not have the ability to generate segmentation output.
-Nonetheless, they still provide quite challenging scenarios that can act as a strong benchmark for the pixel-level MLLMs counterpart.
-In this work, we extend these two benchmarks to augment it with pixel-level annotations and referring expressions that correspond to the object of interest within their VQA task.
-These help to understand if failures occurring in these benchmarks stem from grounding or VQA capabilities.
+Nonetheless, they provide quite challenging scenarios that can act as a strong benchmark for pixel-level MLLMs.
+In this work, we extend these two benchmarks with segmentation annotations and referring expressions that correspond to the object of interest within the VQA task, and propose a paired evaluation metric.
+Moreover, we evaluate these models within a prompt sensitivity evaluation for both language and visual prompts following previous works [22, 4].
+We further introduce novel visual prompt variations specifically designed for the grounding task, pushing the limits of pixel-level MLLMs.
 3 Method and benchmarks
-In this section, we describe our two benchmarks and probing techniques for pixel-level MLLMs and MLLMs that were not trained with pixel-level grounding supervision.
-3.1 Benchmarks
-PixMMVP benchmark: We build upon the recently released MMVP [19] which identified clip blind pairs and used them to build a challenging benchmark with the corresponding questions and choices for 300 images.
-We augment the aforementioned dataset and manually annotate each question with the corresponding object of interest referring expression, e.g.
-an elderly person or the butterfly’s feet.
-There are seven questions only that are not designed to inquire about a specific object in the scene, which are excluded.
-Examples include questions inquiring on the view direction of the camera which are not tied to a specific entity.
-Our manual referring expression annotations are as fine-grained as possible.
-These expressions correspond to what needs to be grounded in the image to answer the question.
+In this section, we describe our two benchmarks and probing techniques for pixel-level MLLMs and MLLMs that were not trained with mask supervision.
+We focus on the referring segmentation as a representative task of visual grounding.
+Furthermore, we use visual question answering as a representative task of their vision-language capabilities, with emphasis on the language.
+3.1 Paired benchmarks for VQA and grounding
+PixMMVP benchmark.
+We build upon the recently released MMVP [24], which identified clip blind pairs and used them to build a challenging benchmark with the corresponding questions and choices for 300 images.
+We manually annotate each question with the corresponding object of interest referring expression.
+There are seven questions only that are not designed to inquire about a specific object in the scene, which are excluded, such as questions inquiring on the view direction of the camera.
+The referring expressions in our dataset correspond to what fine-grained parts need to be grounded in the image to answer the question.
 Afterwards, we manually label these objects of interest with polygonal annotations using the VGG annotator [6].
-PixCV-Bench benchmark: For this benchmark we build upon the 2D component of the recently released CV-Bench [18].
+Hence, we create the first paired benchmark for both VQA and pixel-level visual grounding.
+PixCV-Bench benchmark.
+For this benchmark we build upon the 2D component of the recently released CV-Bench [23].
 We specifically select the 2D component, since they are sourced from segmentation datasets (i.e., ADE20K [35] and COCO [13]), which can be used in our proposed benchmark.
@@
 As such we use GPT-4o to parse the questions and identify the objects of interest automatically, followed by manual inspection and correction.
-Specifically, we collect the classes in each image from the corresponding dataset and construct a list of class choices “1.
-CLS1, 2.
-CLS2, …”.
-Then we prompt GPT-4o with the following, “Provide number only as an answer.
-Identify the objects of interest in the following question: QUESTION ?
-1.
-CLS1, 2.
-CLS2, … ”.
-This provides us with the categories per question that highlights the objects of interest.
-While seemingly these are categorical annotations not referring expressions, certain scenarios in CV-Bench are different.
-Specifically, in the relative positioning task all the questions that include an object highlighted by a red box in the image are annotated with the referring expression, “(annotated by the red box)”, beyond simple categorical annotations.
-Afterwards, we use the selected categories from GPT-4o to retrieve the corresponding segmentation mask/s per image.
-Furthermore, we use a custom annotation tool to manually filter the objects in the question, e.g.
-selecting only the object mask annotated by the red box when referred to it and filtering out the other instances for that same class.
-Another example that needs manual filtration when the class in question is a broader category than what is inquired upon, e.g., “Pendant Lamp” which is under the category of “Lamp” in ADE20K.
-In such a case, we filter out the masks of other types such as “Table Lamp”.
-Moreover, we identify missing annotations in rare occasions that require additional intervention and manually annotate these missing objects.
-We provide the final PixCV-Bench with referring expressions and their segmentation annotations that can be used to evaluate the grounding ability in relation to the original VQA task.
-Appendix A provides visual examples from our benchmarks.
-3.2 A Pixel-level MLLMs study
-We utilize the two proposed benchmarks, PixMMVP and PixCV-Bench, to evaluate how the current trend in pixel-level MLLMs that relies on training with grounding supervision perform in such challenging tasks.
-Furthermore, we inspect the failures of these pixel-level MLLMs and explore simple approaches to pixel-level understanding from MLLMs that overcome the previous shortcomings.
-Pixel-level MLLMs shortcomings.
-We highlight the failures for the current state-of-the-art pixel-level MLLMs through three probing techniques.
-First, we highlight the degraded performance in VQA from most of the pixel-level MLLMs that are trained with pixel-level grounding supervision.
-We use for that the following prompt, “IMGQUESTION?
-OPTION1 OPTION2…”, as shown in Figure 2.
-Certain pixel-level MLLMs tend to answer the aforementioned question while outputting a corresponding segmentation mask/s for the objects of interest.
-Notably, the worst two models in this task, LISA [9] and GLAMM [16], are not able to provide an answer and rather refer to a segmentation mask.
-On the other hand, OMG-Llava [27] shows better ability in VQA.
-The second shortcoming we discuss is their degraded ability to visually ground objects.
-Surprisingly, although they were trained with pixel-level grounding supervision, not all of these models show superior grounding performance.
-Figure 3 shows the second prompt to generate a segmentation mask for the ground-truth referring expression.
-The purpose of this probing is to understand whether the failure in these models is purely on the VQA task, or its inability to ground the objects of interest in the corresponding question or both.
-Figure 3 shows the worst two models in this aspect, which are GLAMM, the region captioning variant, and Llava-G.
-Both fail to segment the specific object in question, while OMG-Llava shows better performance.
-Third, we highlight another shortcoming, where these MLLMs exhibit degraded ability to follow instructions.
+This provides us with the categories per question that highlight the objects of interest.
+While seemingly these are categorical annotations, not referring expressions, certain scenarios in CV-Bench are different.
+Specifically, in the relative positioning task, all the questions that include an object highlighted by a red box in the image are annotated with the referring expression, “(annotated by the red box)”, beyond simple categories.
+Afterwards, we use the expressions from GPT-4o to retrieve the corresponding segmentation mask per image.
+Furthermore, we use a custom annotation tool to manually filter the objects in question, e.g.
+selecting only the object mask annotated by the red box and filtering out other instances.
+Another example that needs manual filtration is when the class in question is a broader category than what is inquired about.
+Moreover, we identify missing annotations and manually annotate these missing objects.
+We provide the final paired PixCV-Bench with referring expressions, their segmentation annotations, visual questions and corresponding answers.
+The supplementary provides visual examples from our benchmarks.
+3.2 Pixel-level MLLMs study
+We utilize the two proposed benchmarks, PixMMVP and PixCV-Bench, and perform an additional prompt sensitivity analysis to evaluate and inspect the failures of these pixel-level MLLMs.
+The goal is to push these models in the direction of learning visual grounding abilities on the pixel-level without sacrificing the language abilities that are integral to grounding.
+Pixel-level MLLMs failures.
+We highlight the failures of the current state-of-the-art pixel-level MLLMs through three probing techniques.
+First, we highlight the degraded performance in VQA from most of these MLLMs that are trained with mask supervision.
+We use the following prompt, “IMGQUESTION?
+OPTION1 OPTION2…”, as shown in Fig.
+2a.
+Notably, the worst models in this task, e.g., GLAMM [20], are not able to provide an answer and rather refer to a segmentation mask.
+On the other hand, OMG-LLaVA [34] tries to tackle the question, even if incorrectly.
+The second failure we discuss is that these MLLMs exhibit a degraded ability to follow instructions.
 In order to probe this, we use the following prompt: “IMGQUESTION?
-a.OPTION1 b.OPTION2… Answer with the option’s letter from the given.” Figure 4 shows an example with the answers from the worst two models in this aspect which are LISA [9] and Llava-G [26].
-Both are incapable of following the instruction, yet Llava-G tries to tackle the question unlike LISA.
-On the other hand, OMG-Llava shows better ability to follow the instruction and answer the question.
-Baselines and upper bounds.
-In addition to evaluating state-of-the-art pixel-level MLLMs, we propose two baselines and one upper bound.
-The first of which is inspired by a concurrent work [2] that identified the emergent grounding in multi-modal large language models without the need for any pixel-level grounding supervision.
+a.OPTION1 b.OPTION2 … Answer with the option’s letter from the given.” Figure 2b shows an example from one of the worst models in this aspect which is LLaVA-G [33] that is incapable of following the instruction, yet tries to tackle the question in free-form.
+On the other hand, OMG-LLaVA shows better ability to follow the instructions and answer the question.
+Third, we highlight their degraded ability to visually ground objects.
+Surprisingly, although they were trained with mask supervision for grounding, not all of these models show superior grounding performance on our benchmarks.
+Figure 2c shows the prompt to generate a segmentation mask for the queried referring expression.
+The purpose of this probing is to understand whether the failure in these models is purely in its language capabilities, or its inability to ground the objects of interest in the corresponding question or both.
+Figure 2c shows LLaVA-G and OMG-LLaVA, where the latter shows better performance.
+Prompt sensitivity evaluation.
+We propose language and visual prompts variations tailored to our paired benchmark to measure the sensitivity of both pixel-level and standard MLLMs.
+We argue that language is an integral component in the visual grounding task and, consequently, better language abilities in MLLMs can entail better grounding.
+Figure 3 shows a Qwen2.5-VL failure from a language prompt variation, where both ask the model to locate the dog’s face, yet the model fails in the second prompt.
+This motivates our paired evaluation study, which is augmented with a prompt sensitivity evaluation.
+We conduct the prompt sensitivity evaluation on PixMMVP, since the dataset is considered out-of-distribution with respect to what the majority of MLLMs were instruction-tuned on, unlike PixCV-Bench.
+For the language variations, we follow previous works [4] that rely on three variants.
+These include: (i) spelling errors from omission, transposition, and addition of letters that are randomly generated across eight random locations in the text of the question, the choices and instructions, (ii) random selection of the template for the question and choices and (iii) paraphrasing of the question using GPT-4o.
+For the visual prompts variations, we propose a novel method that is tailored for visual grounding, where we propose guided random cropping and masking.
+In the former, the ground-truth mask is used to crop more than 50% of the referred object, while the latter uses a black rectangle to randomly mask out more than 50% of it.
+Additional details are included in the supplementary.
+| Image | Referring Expression | Concept Category | Noun Phrase | Output |
+| 1 | the butterfly’s wings | Color & Appearance | orange wings | In the image, there is a butterfly with orange wings. |
+| 3 | the flame of the match | Location & Position | the top | The flame of the match is located at the top of the image, surrounded by darkness. |
+| 6 | the dog’s face | Color & Appearance | a black and white dog | The dog’s face in the scene is a black and white dog with a black nose. |
+| 161 | the minute hand of the clock | Location & Position | the 12 o’clock position | The minute hand of the clock in the scene is located at the 12 o’clock position. |
+1
+3
+6
+161
+3.3 Interpretability mechanism
+In addition to evaluating state-of-the-art pixel-level MLLMs, we propose a novel interpretability mechanism to study when grounding emerges in MLLMs with respect to the output tokens.
+We are inspired by a concurrent work [3] that identified the emergent pixel-level grounding in MLLMs without the need for any mask supervision.
 Specifically, we use their attend and segment meta architecture as one of our baselines.
-However, we are the first to discuss when does such grounding emerge in these models.
-We identify an interesting connection between the identified output tokens and the output grounding from the attention maps that gives insights on how these models reason.
-The attend and segment meta-architecture extracts the raw attention map for the output token, , where are the number of layers and heads, resp.
-Then, are the number of input language tokens before and after the visual tokens respectively, while are the height and width of the input image.
-Only the attention corresponding to the visual tokens of length are used, and these attention maps are averaged across the layers and heads, resulting in .
+However, we are the first to discuss when such grounding emerges in these models.
+We identify an interesting connection between the identified output tokens and the output grounding from the attention maps that gives insights into how these models reason.
+We extract the raw attention map for the output token, , where are the number of layers and heads, resp.
+Then, are the number of input language tokens before and after the visual tokens, respectively, while are the height and width of the input image features.
+Only the attention corresponding to the visual tokens of length is used, and these attention maps are averaged across the layers and heads, resulting in .
 This is further normalized across all the output, for output tokens.
-The attend and segment depends on using spaCy natural language processing tool [6] to identify the noun phrases and associate them to the ground-truth referring expressions.
-Thus, the spaCy embeddings closest to the ground-truth expression are used in the mask selection.
+The attend and segment baseline depends on using the spaCy natural language processing tool [8] to identify the noun phrases and associate them with the queried referring expressions.
+Thus, the spaCy embeddings closest to the queried expression are used in the token selection and their respective attention maps, , are averaged.
 This is followed by extracting the maximum attention point to feed into SAM [11] as a point prompt.
-For our baseline and upper bound, we build upon the previous pipeline and build an oracle upper bound and an automatic baseline.
-We introduce two main modifications to account for our observation that the correct grounding can occur with different output tokens describing the object not necessarily aligning with the exact ground-truth expression.
-The first modification is to go through all the potential output tokens without relying on spaCy embeddings.
-In the oracle we rely on the ground-truth mask to select the correct token and its corresponding attention map with highest intersection over union as an upper bound.
-The automatic baseline uses a simple but powerful mechanism where we overlay the predicted masks on the original image to highlight the potential object of interest.
-This is followed by feeding these images to a multi-modal LLM inquiring on which is best in highlighting this object.
-Specifically, we use the following prompt “Select the image that has EXPR best highlighted in red color than the others?
-Answer with a number from 1 to N and mention the number only.
-IMG”, where EXPR and IMG are the ground-truth expression and the image tokens respectively.
-In our automatic baseline we rely on GPT-4o for the mask selection.
-The second modification, since SAM has a good understanding of point prompting ambiguity, we process three potential output masks for each prompt instead of one only.
-This enables us to utilize the power of SAM in identifying fine-grained objects and referring expressions that tends to surpass what other MLLMs do, even those trained with pixel-level grounding supervision.
-| Method | PixGr. | PixMMVP | PixCV-Bench | ||||||||
-| Llava 1.5 (7B) [11] | ✗ | 27.3 | 28.0 | - | - | - | 17.4 | 60.3 | - | - | - |
-| Llava 1.5 (13B) [11] | ✗ | 39.3 | 30 | - | - | - | 14.5 | 61.4 | - | - | - |
-| Cambrian (8B)* [18] | ✗ | 52.0 | 52.0 | - | - | - | 62.2 | 72.2 | - | - | - |
-| OMG Llava (7B)** [27] | ✓ | 12.0 | 12.0 | 17.8 | 38.0 | 18.2 | 12.0 | 42.1 | - | 50.5 | 45.9 |
-| GLAMM (7B) [16] | ✓ | 1.3 | 2.7 | 31.5 | 47.4 | 5.1 | - | - | 30.2 | 51.9 | - |
-| GLAMM - RegCap (7B) [16] | ✓ | 12.7 | 6.7 | 14.5 | 18.6 | 15.1 | 27.8 | 54.4 | 3.6 | 7.4 | 13.0 |
-| LISA (7B) [9] | ✓ | 7.3 | - | 18.1 | 42.9 | 12.5 | 3.7 | - | 16.8 | 48.1 | 6.7 |
-| Llava-G (7B) [26] | ✓ | 9.3 | - | 17.8 | 13.5 | 12.2 | 14.1 | 4.4 | 1.7 | 17.6 | 15.8 |
-| Llava 1.5 (7B) + (a+s) [2] | ✗ | 27.3 | 28.0 | 11.1 | 11.2 | 16.0 | 17.4 | 60.3 | 5.2 | 15.7 | 24.9 |
-| Llava 1.5 (13B) + (a+s) [2] | ✗ | 39.3 | 30 | 9.8 | 11.4 | 17.7 | 14.5 | 61.4 | 4.7 | 14.9 | 24.0 |
-| Cambrian (8B)* + (a+s) [2] | ✗ | 52.0 | 52.0 | 14.3 | 15.1 | 23.4 | 62.2 | 72.2 | 18.6 | 15.9 | 29.6 |
-| PixFoundation (7B) (Ours) | ✗ | 27.3 | 28.0 | 18.8 | 25.9 | 26.9 | 17.4 | 60.3 | 5.4 | 28.5 | 38.7 |
-| PixFoundation (13B) (Ours) | ✗ | 39.3 | 30 | 16.9 | 25.0 | 30.6 | 14.5 | 61.4 | 4.8 | 27.6 | 38.1 |
-| PixFoundation (8B)* (Ours) | ✗ | 52.0 | 52.0 | 29.6 | 30.3 | 38.3 | 62.2 | 72.2 | 23.9 | 33.1 | 45.4 |
-| Upper Bound - Oracle Selection | |||||||||||
-| PixFoundation (7B) (Ours) | ✗ | 27.3 | 28.0 | 26.1 | 38.0 | 32.2 | 17.4 | 60.3 | 6.3 | 49.7 | 54.5 |
-| PixFoundation (13B) (Ours) | ✗ | 39.3 | 30 | 23.6 | 38.2 | 38.7 | 14.5 | 61.4 | 5.3 | 50.6 | 55.5 |
-| PixFoundation (8B)* (Ours) | ✗ | 52.0 | 52.0 | 52.0 | 56.1 | 54.0 | 62.2 | 72.2 | 54.3 | 64.4 | 68.1 |
 the dorsal fin of the animal
-| Image | Prompt | Concept Category | Noun Phrase | Output |
-| 1 | Identify the butterfly’s wings in the scene. | Color & Appearance | orange wings | In the image, there is a butterfly with orange wings. |
-| 3 | Identify the flame of the match in the scene. | Location & Position | the top | The flame of the match is located at the top of the image, surrounded by darkness. |
-| 6 | Identify the dog’s face in the scene. | Color & Appearance | a black and white dog | The dog’s face in the scene is a black and white dog with a black nose. |
-| 161 | Identify the minute hand of the clock in the scene. | Location & Position | the 12 o’clock position | The minute hand of the clock in the scene is located at the 12 o’clock position. |
-[5pt]
-1
-[5pt]
-3
-[5pt]
-6
-[5pt]
-161
+For our interpretability mechanism, we build upon the previous pipeline and provide an oracle upper bound and an automatic method, which we call PixFoundation.
+We propose an additional mask selection by using MLLM as a judge on the different segmentation masks corresponding to the output tokens, to select the best mask and its respective noun phrase.
+Specifically, for each noun phrase in the output, we average the attention maps, , for each respective token, .
+Afterwards, we extract the maximum attention point, prompt SAM with it, highlight the generated mask in the image and ask the MLLM to select the highlighted image that best describes the queried referring expression from all the masks and their respective noun phrases.
+Furthermore, we consider that the point prompt ambiguity is crucial in identifying fine-grained details; as such, we allow for our mask selection to consider three masks from SAM for each point and its respective noun phrase.
+The automatic method can be used with either closed-source or open-source MLLMs without requiring groundtruth segmentation.
+It can be used as an interpretability tool to have a better understanding of where MLLMs look and when the correct grounding occurs with respect to the output tokens.
+In the oracle, we rely on the ground-truth mask to select the correct token and its corresponding segmentation with the highest intersection over union as an upper bound.
+More details are provided in the supplementary.
+Figure 4 shows qualitative examples from our interpretability tool, the oracle variant, applied to LLaVA 1.5, where the best masks and their respective noun phrases do not necessarily correspond to the queried referring expression.
+We also study the concept categories for the noun phrases where the best grounding occurs, relying on our interpretability tool, refer to Sec.
+4.3.
+Figure 5 shows how the output from our interpretability mechanism and the oracle variant is better capable of identifying the fine-grained detail, “dorsal fin”, from LLaVA 1.5 without being trained with mask supervision, unlike other pixel-level MLLMs, while still maintaining its language capabilities.
+Considering that pixel-level MLLMs rely on standard ones as their base MLLM, our benchmarking and interpretability study questions whether there could be better pixel-level MLLMs built without the loss of their original language and grounding capabilities.
+| Method | Venue | Grounding Output | PixMMVP | PixMMVP Prompt Sensitivity | |||||||
+| w/ GPT | |||||||||||
+| Pixel-level MLLMs w/ Mask Output | |||||||||||
+| OMG LLaVA (7B) | NeurIPS 2024 | Mask | 12.0 | 12.0 | 17.8 (13.8) | 38.0 (38.8) | 18.2 (18.3) | 5.9 | 11.8 | 29.7 | 9.8 |
+| GLAMM (7B) | CVPR 2024 | Mask | 1.3 | 2.7 | 31.5 (30.6) | 47.4 (48.3) | 5.1 (5.1) | - | - | 41.9 | - |
+| LISA (7B) | CVPR 2024 | Mask | 7.3 | - | 18.1 (14.1) | 42.9 (42.9) | 12.5 (12.5) | - | - | 35.0 | - |
+| LLaVA-G (7B) | ECCV 2024 | Mask | 9.3 | - | 19.2 (15.2) | 28.8 (29.1) | 14.0 (14.1) | - | - | 26.8 | - |
+| RGA (7B) | ICCV 2025 | Mask | 5.3 | 51.3 | - | 52.7 (54.6) | 52.0 (52.9) | 23.2 | 28.5 | 45.8 | 35.1 |
+| Standard MLLMs Extended for Mask Output | |||||||||||
+| LLaVA 1.5 (7B) + SAM | baseline (ours) | Box2Mask | 27.3 | 28.0 | - | 38.2 (40.1) | 32.3 (33.0) | 18.4 | 18.7 | 21.9 | 20.2 |
+| LLaVA 1.5 (13B) + SAM | baseline (ours) | Box2Mask | 39.3 | 30 | - | 45.5 (47.7) | 42.2 (43.1) | 24.4 | 23.7 | 24.9 | 24.6 |
+| Cambrian (8B) + SAM | baseline (ours) | Box2Mask | 52.0 | 52.0 | - | 39.9 (41.8) | 45.2 (46.3) | - | 44.2 | 28.4 | 34.6 |
+| Qwen2.5-VL (7B) + SAM | baseline (ours) | Point2Mask | 54.7 | 38.0 | - | 37.6 (36.3) | 44.6 (43.6) | 29.6 | 33.8 | 24.8 | 28.6 |
+| Qwen2.5-VL (7B) + SAM | baseline (ours) | Box2Mask | 54.7 | 38.0 | - | 33.2 (31.0) | 41.3 (39.6) | 29.6 | 33.8 | 26.0 | 29.4 |
+| LLaVA 1.5 (7B) + (a+s) | baseline (Arxiv 2025) | Point2Mask | 27.3 | 28.0 | 11.8 (7.8) | 18.1 (14.1) | 22.0 (18.8) | 18.4 | 18.7 | 16.5 | 17.5 |
+| LLaVA 1.5 (13B) + (a+s) | baseline (Arxiv 2025) | Point2Mask | 39.3 | 30 | 12.1 (8.5) | 17.9 (13.9) | 24.6 (20.5) | 24.4 | 23.7 | 16.3 | 19.5 |
+| Cambrian (8B) + (a+s) | baseline (Arxiv 2025) | Point2Mask | 52.0 | 52.0 | 19.5 (15.5) | 16.7 (13.3) | 28.4 (23.9) | - | 44.2 | 18.8 | 26.4 |
+| LLaVA 1.5 (7B) + PixFoundation | interp. tool (ours) | Point2Mask | 27.3 | 28.0 | 31.2 (29.2) | 42.8 (42.1) | 33.9 (33.6) | 18.4 | 18.7 | 37.9 | 25.0 |
+| LLaVA 1.5 (13B) + PixFoundation | interp. tool (ours) | Point2Mask | 39.3 | 30 | 27.1 (26.0) | 41.3 (40.1) | 40.3 (39.7) | 24.4 | 23.7 | 37.5 | 29.6 |
+| Cambrian (8B) + PixFoundation | interp. tool (ours) | Point2Mask | 52.0 | 52.0 | 47.0 (44.7) | 41.8 (41.4) | 49.4 (48.1) | - | 44.2 | 33.6 | 38.2 |
+| Qwen2.5-VL (7B) + PixFoundation | interp. tool (ours) | Point2Mask | 54.7 | 38.0 | - | 46.5 (44.2) | 50.3 (48.9) | 29.6 | 33.8 | 31.6 | 32.7 |
+| LLaVA 1.5 (7B) + PixFoundation | upper bound (ours) | Point2Mask | 27.3 | 28.0 | 37.8 (34.8) | 55.3 (53.1) | 37.2 (36.7) | 18.4 | 18.7 | 56.4 | 28.1 |
+| LLaVA 1.5 (13B) + PixFoundation | upper bound (ours) | Point2Mask | 39.3 | 30 | 36.8 (33.7) | 55.2 (53.0) | 45.9 (45.1) | 24.4 | 23.7 | 55.2 | 33.8 |
+| Cambrian (8B) + PixFoundation | upper bound (ours) | Point2Mask | 52.0 | 52.0 | 68.7 (67.1) | 72.1 (70.8) | 60.4 (60.0) | - | 44.2 | 70.9 | 54.5 |
+| Qwen2.5-VL (7B) + PixFoundation | upper bound (ours) | Point2Mask | 54.7 | 38.0 | - | 55.0 (52.8) | 54.8 (53.7) | 29.6 | 33.8 | 38.9 | 36.2 |
+| Method | Grounding Output | PixCV-Bench | |||
+| Pixel-level MLLMs w/ Mask Output | |||||
+| OMG LLaVA (7B) | Mask | 12.0 | 42.1 | 50.5 | 45.9 |
+| GLAMM (7B) | Mask | - | - | 51.9 | - |
+| LISA (7B) | Mask | 3.7 | - | 48.1 | 6.7 |
+| LLaVA-G (7B) | Mask | 14.1 | 4.4 | 17.6 | 15.8 |
+| RGA (7B) | Mask | - | 72.6 | 51.1 | 56.0 |
+| Standard MLLMs Extended for Mask Output | |||||
+| LLaVA 1.5 (7B) + SAM | Box2Mask | 17.4 | 60.3 | 29.8 | 39.9 |
+| LLaVA 1.5 (13B) + SAM | Box2Mask | 14.5 | 61.4 | 32.9 | 42.8 |
+| Cambrian (8B) + SAM | Box2Mask | 62.2 | 72.2 | 34.2 | 46.4 |
+| Qwen2.5-VL (7B) + SAM | Point2Mask | 48.9 | 60.8 | 33.7 | 43.4 |
+| Qwen2.5-VL (7B) + SAM | Box2Mask | 48.9 | 60.8 | 48.9 | 54.2 |
+| LLaVA 1.5 (7B) + (a+s) | Point2Mask | 17.4 | 60.3 | 15.7 | 24.9 |
+| LLaVA 1.5 (13B) + (a+s) | Point2Mask | 14.5 | 61.4 | 14.9 | 24.0 |
+| Cambrian (8B) + (a+s) | Point2Mask | 62.2 | 72.2 | 15.9 | 26.1 |
+| LLaVA 1.5 (7B) + PixFoundation | Point2Mask | 17.4 | 60.3 | 34.1 | 43.6 |
+| LLaVA 1.5 (13B) + PixFoundation | Point2Mask | 14.5 | 61.4 | 33.9 | 43.7 |
+| Cambrian (8B) + PixFoundation | Point2Mask | 62.2 | 72.2 | 38.4 | 50.1 |
+| LLaVA 1.5 (7B) + PixFoundation | Point2Mask | 17.4 | 60.3 | 49.7 | 54.5 |
+| LLaVA 1.5 (13B) + PixFoundation | Point2Mask | 14.5 | 61.4 | 50.6 | 55.5 |
+| Cambrian (8B) + PixFoundation | Point2Mask | 62.2 | 72.2 | 64.4 | 68.1 |
+| Qwen2.5-VL (7B) + PixFoundation | Point2Mask | 48.9 | 60.8 | 43.6 | 50.8 |
 4 Experiments
@@
 PixMMVP is composed of 300 images paired with questions, choices, referring expressions and segmentation masks, while PixCV-Bench has 1,438 images with their corresponding annotations similarly.
-On each benchmark we evaluate the VQA and visual grounding capabilities following three probing techniques and reporting their metrics.
+On each benchmark, we evaluate the VQA and referring segmentation following three probing techniques and report their metrics.
 The first probing is to evaluate the VQA ability, where the accuracy is computed using GPT-4o following [24] as, .
-If the model generates a segmentation without explicitly asking it to, it is evaluated with respect to the ground-truth referring segmentation in terms of mean intersection over union as, .
-The second probing prompts the model to identify the referred expression then evaluates the mean intersection over union reported as, .
-The third probing following [18] instructs the model to generate a single option letter and evaluate the accuracy directly without GPT-4o, reported as, .
-There is a need for the first probing since some of the recent pixel-level MLLMs face challenges in following instructions.
-We evaluate the score of each model, , which is the harmonic mean across the maximum of both pixel-level visual grounding and VQA,
-We mainly focus on evaluating four state-of-the-art pixel-level MLLMs; LISA [9], GLAMM [16], OMG-Llava [27] and Llava-G [26].
-For GLAMM we use two variants; the original model (GLAMM) and the one fine-tuned for region-level captioning, (GLAMM-RegCap).
-For details on the models weights, refer to App.
-A.
-Baselines and upper bound implementation details.
-We evaluate: (i) the attend and segment (a+s), (ii) the oracle selection relying on the highest intersection over union in selecting the correctly predicted masks (PixFoundation), and (iii) the automatic selection, (PixFoundation).
-These are implemented on top of three base MLLMs which are, Llava 1.5 (7B, 13B) [12] and Cambrian-1(8B) [18].
-The automatic selection is implemented using GPT-4o, refer to App.
-A for more details.
+If the model generates a segmentation without explicitly asking it to, it is evaluated with respect to the ground-truth referring segmentation in terms of mean intersection over union as .
+The second probing prompts the model to identify the referred expression and evaluates the mean intersection over union reported as .
+The third probing following [23] instructs the model to generate an option letter and evaluate the accuracy directly without GPT-4o, reported as .
+We evaluate the score of each model, , which is the harmonic mean across both referring segmentation and VQA,
+For the language prompt sensitivity evaluation, we randomly generate 10 variations for each of the three language variants, and for the visual one, we generate four for each of the two variants.
+As such, we have in total 30 variations for the VQA and eight for referring segmentation; for simplicity, we report the mean of the metric used across all the variations.
+For the VQA accuracy, we follow the third probing, where accuracy is evaluated using the generated option’s letter with simple postprocessing on the model’s output.
+To avoid observations from previous work [9] that prompt sensitivity failures can stem from the evaluation and post-processing itself, we evaluate with GPT-4o the same output and report both, as and w/ GPT, respectively.
+We provide in the supplementary materials a language prompt sensitivity evaluation for the grounding aside from the VQA.
+Compared methods.
+We focus on evaluating five state-of-the-art pixel-level MLLMs; LISA [12], GLAMM [20], OMG-LLaVA [34], LLaVA-G [33] and RGA [26].
+These were either equipped with a SAM or SAM 2 decoder or with other powerful decoders, e.g., OMG-Seg.
+Some of these were trained with the SAM dataset, SA-1B, extended to the visual grounding [20] or other large-scale segmentation data.
+Furthermore, we evaluate the attend and segment (a+s) [3], the oracle relying on the highest intersection over union in mask selection (PixFoundation), the automatic selection (PixFoundation) and a simple baseline that uses the output bounding box/point to prompt SAM without any selection.
+These are implemented on top of four base MLLMs, which are LLaVA 1.5 (7B, 13B) [14], Cambrian (8B) [23] and Qwen2.5-VL [1].
+The automatic selection is implemented using GPT-5.1.
+Additional details are in the supplementary.
 4.2 Are the current pixel-level MLLMs heading in the right direction?
-In order to answer this, we evaluate each of these pixel-level MLLMs VQA capabilities in challenging tasks.
-Additionally, we evaluate their ability to visually ground the objects of interest in these questions.
-Table 1 shows the results on the challenging PixMMVP and PixCV-Bench.
-From the accuracy of VQA, MLLMs that are not trained with pixel-level grounding surpass their pixel-level counterpart with up to 16%.
-The best in pixel-level MLLMs score in this aspect is GLAMM-RegCap [27] yet it has degraded ability to generate segmentation.
-On the other hand, when looking at pixel-level visual grounding we find the best model, GLAMM [16], has a weak ability in VQA or following instructions.
-Moreover, it shows LISA and Llava-G are mostly incapable of following the instruction to output the option letter reported in, .
-OMG-Llava strikes the right balance in both VQA and pixel-level grounding with the highest score, , within pixel-level MLLMs.
-However, looking at the bottom three rows, the oracle, PixFoundation, confirms that MLLMs that were never trained with pixel-level grounding have the correct grounding within their learned attention maps yielding higher score, , than OMG-Llava.
-Qualitative comparison is provided in Fig.
-5 and additional examples are in App.
-B.
-Additionally, the automatic baseline outperforms OMG-Llava in PixMMVP with a considerable margin and is on-par to it on PixCV-Bench.
-Furthermore, the attend and segment baseline [2] lags behind our automatic method with more than 10% in both datasets.
-Refer to App.
-C for a failure case analysis.
-Finally, we evaluate whether the failures of these MLLMs occur in visual grounding, VQA or both.
-Figure 6 shows the frequency of failures per category, where the majority stem from failures in both especially in the pixel-level MLLMs.
-While the vanilla MLLMs perform better in the VQA and tend to fail more in the grounding task.
+In order to answer this, we evaluate each of these pixel-level MLLMs’ capabilities in VQA and referring segmentation in challenging tasks.
+Table 1 shows the results on the challenging PixMMVP and the respective prompt sensitivity ones.
+From the accuracy of VQA, standard MLLMs that are not trained for mask output surpass their pixel-level counterpart with 3-14%.
+This is especially evident in the language prompt sensitivity, where RGA, the best pixel-level MLLM, drops VQA accuracy to 28.5%, which is worse than other standard MLLMs, i.e., Cambrian (8B) and Qwen2.5-VL (7B) at 44.2% and 33.8%, respectively.
+Furthermore, RGA shows severe failure in , which is attributed to its sensitivity to the VQA prompt, as it can only answer the multiple-choice question when instructed to generate a letter, while exhibiting limited ability to generate free-form answers.
+As for the referring segmentation, some of these pixel-level MLLMs, e.g., OMG-LLaVA and LLaVA-G, underperform the baselines.
+Comparing PixFoundation and the baseline (a+s) clearly shows that our interpretability tool is better because of the understanding that the grounding does not necessarily coincide with the output noun phrase closest to the referring expression, unlike the (a+s) baseline.
+This is further confirmed in the point accuracy results in Table 3.
+When looking at the overall score, , in PixMMVP, our concurrent work (RGA) shows strong performance that is better than its respective model Qwen2.5-VL, yet in the prompt sensitivity analysis, it shows lower performance than Cambrian with PixFoundation.
+The second-best pixel-level MLLM, OMG-LLaVA, which was prior to the release of our benchmark, consistently lags behind our baselines.
+Various factors can impact the language and visual abilities of these MLLMs, especially in visual grounding and VQA.
+These results could be attributed to training recipes, design choices or datasets; this is evident with RGA performance surpassing other pixel-level MLLMs, which rely on mixing both QA and segmentation datasets during training and expanding them with video tasks.
+However, even with RGA, its VQA abilities downgraded drastically in the language prompt sensitivity with a factor of 1.8, while the standard MLLM Cambrian degraded with a factor of 1.2.
+Unlike RGA, Qwen2.5-VL and Cambrian generate grounding output directly in language.
+There could be potential opportunities to improve the language capabilities through exploring these design choices further in the segmentation domain.
+Table 2 shows PixCV-Bench results.
+Our concurrent work, RGA, shows strong performance, followed closely by Qwen2.5-VL.
+It also shows that Point2Mask is worse than Box2Mask in our simple baselines in PixCVBench, which could be due to the referring expressions inquiring about multiple objects/instances more than in PixMMVP due to the nature of the questions, e.g., relative positioning.
+Furthermore, Point2Mask shows higher sensitivity to visual prompt variations than Box2Mask.
 Summary.
-In summary, pixel-level grounding supervision degrades MLLMs ability in VQA and sometimes even their generalization in grounding.
-We show that MLLMs trained with pixel-level supervision lag behind vanilla MLLMs using simple mechanisms to extract grounding, and the oracle indicates there is an opportunity to improve this.
-Moreover, we show that grounding might not coincide with the noun phrase most similar to the referred expression, where our oracle upper bound and automatic baseline are both surpassing the attend and segment.
+In summary, the state-of-the-art pixel-level MLLMs prior to the release of our benchmark have shown lower performance than simple baselines, where they specifically lag in their language abilities and visual question answering.
+Prior to our work, pixel-level MLLMs were mostly ignoring the language abilities, yet we show various examples that emphasize the role of language in visual grounding.
+Moreover, we show that visual grounding might not coincide with the output noun phrase most similar to the referred expression, where our oracle upper bound and automatic method both surpass the attend and segment.
+Our interpretability tool enabled such an important insight to help understand these MLLMs better.
+| Base MLLM | PixFoundation | a+s | PixFoundation |
+| LLaVA 1.5 (7B) | 58.7 | 24.3 | 51.4 |
+| LLaVA 1.5 (13B) | 66.0 | 28.3 | 58.7 |
 4.3 When does grounding emerge in MLLMs?
 When - location.
-Taking into account the powerful performance of the oracle upper bound, it begs the important question of when grounding emerges.
-We start by looking at when it emerges in terms of the location within the output text.
-We analyze the word/phrase location with respect to the full output text in terms of a percentage from its total length, (i.e., 0% means the beginning of the text).
+Taking into account the powerful performance of the oracle upper bound, it begs the question of when grounding emerges.
+We start by looking at when it emerges in terms of the location.
+We analyze the word/phrase location with respect to the full output text in terms of a percentage of its total length (i.e., 0% means the beginning of the text).
 Accordingly, Fig.
-7(a) shows the location percentages histogram, binned at 10%, for the three base MLLMs reporting the oracle selection and evaluating on PixMMVP benchmark using the second probing.
-In the Llava 1.5 variants, the highest grounding emerges at the last 40% of the output text, while for Cambrian it emerges at the last 60%.
+6(a) shows the location percentages histogram, binned at 10%, for the three base MLLMs reporting the oracle selection and evaluating on PixMMVP benchmark using the second probing.
+In the LLaVA 1.5 variants, the highest grounding is at the last 40%, while for Cambrian it is at the last 60%.
 When - concept.
-For the second analysis we look into the concept category that the correct output word/phrase corresponds to.
-The previous assumption in other works is that grounding emerges in the exact noun/noun phrase of the object of interest.
+For the second analysis, we look into the concept category that the correct output word/phrase corresponds to.
+The previous assumption, in other works is that grounding emerges in the exact noun/noun phrase of the object of interest.
 Except our analysis confirms that this is not necessarily the case.
-We take the correct noun/noun phrase where the grounding emerges based on the oracle from all the three variants, then we pass it to GPT-4o to request a grouping of these concepts.
-It resulted into six main groups, which are: (i) Color and appearance, (ii) location and position, (iii) object parts, (iv) context and setting, (v) objects and entities, and (vi) State.
-We then prompt for each of the noun/noun phrase, GPT-4o, to categorize it within these six categories.
+We take the correct noun/noun phrase where the grounding emerges based on the oracle from all three variants, then we pass it to GPT-4o to request a grouping of these concepts.
+It results in six main groups, which are: (i) color and appearance, (ii) location and position, (iii) object parts, (iv) context and setting, (v) objects and entities, and (vi) State.
+We then prompt for each of the noun/noun phrases, GPT-4o, to categorize it within these six categories.
 The histogram of the occurrences of these concept categories is shown in Fig.
-7(b).
-It clearly conveys that in certain scenarios the correct output when grounding emerges can be describing the position or the color of the object of interest not necessarily the exact ground-truth referring expression.
-Fig.
-8 shows qualitative examples of these scenarios from PixMMVP using the second probing.
-Additional results are provided in App.
-B.
-Random vs.
-best.
-All the results of our baselines and our findings hinge on the fact that we are using the maximum attention per output noun phrase to prompt SAM for the segmentation mask.
-Nonetheless, as a lower bound analysis, we evaluate the performance if we use a random point as prompt to SAM instead.
+6(b).
+It conveys that in certain scenarios, the correct output when grounding emerges can be describing the position or the color of the object.
+Figure 4 shows qualitative examples of these scenarios.
+Summary.
+In summary, we found that emergent pixel-level grounding might not coincide with the input referring expression.
+We show that grounding in MLLMs can emerge in the noun phrase that corresponds to color, position or other characteristics of the object of interest.
+5 Conclusion
+We propose two benchmarks showing that pixel-level MLLMs degrade the ability in VQA and even grounding of fine-grained objects.
+We propose language and visual prompt variations tailored for visual grounding.
+Additionally, we provide an interpretability mechanism that showed an interesting insight on when visual grounding occurs in MLLMs.
+Our paired benchmarks and evaluation pave the road towards better and interpretable pixel-level MLLMs.
+Appendix A Additional implementation details
+In this section, we cover additional details about our proposed datasets, the implementation of the evaluation setup, prompt sensitivity evaluation and baselines.
+Datasets.
+Our proposed datasets, PixMMVP and PixCV-Bench, are composed of referring expressions describing the object of interest in the respective question and its segmentation mask.
+We show in Fig.
+8 examples of these ground-truth annotations for both datasets.
+It shows the challenging scenarios in pixel-level visual grounding, which is strongly tied to the visual question answering task, since an integral part of answering these questions requires the grounding of the object/s of interest.
+Models.
+We also detail the model checkpoints we use for the four pixel-level MLLMs and their variants, retrieved from HuggingFace [28] in Table 4.
+These also include the model checkpoints used for the base MLLMs that were not trained with mask supervision.
+Furthermore, we provide details on the oracle selection mechanism.
+We discard the cases where the ground-truth segmentation is all background in the oracle selection in the when analysis and the point accuracy evaluation, since there is no ground-truth grounding to evaluate against.
+While in the quantitative and qualitative evaluation, we use an empty mask for the oracle.
+For the automatic variant, we automatically identify the images that do not have the referring expression, and correspondingly output an empty mask.
+Specifically, we prompt the models with the question: “Does this image have EXP?
+Answer with Yes/No only.” Similarly, for the attend and segment architecture (a+s), we follow their approach in filtering out the selected noun phrases that have a similarity less than 0.7 in the SpaCy embeddings and instead output an empty mask.
+This pre-processing is only needed in PixMMVP, while on PixCV-Bench we do not need such filtration across the (a+s), automatic or oracle variants.
+In the oracle upper bound, whenever the model is to be evaluated in a multiple object scenario, we take all the possible pairs of the masks and select the best pair based on the highest intersection over union.
+Additionally, we provide details on the SAM model that is used in our baselines and interpretability tool, where we use the ViT-H variant.
+For the automatic and the oracle variants of PixFoundation with Qwen2.5-VL, we use the output point generated directly from the model instead of mining the attention maps, where we provide these results for reference.
+Nonetheless, we maintain the mask selection among the three masks corresponding to each point.
+Finally, we provide an illustrative example of our automatic selection mechanism with the corresponding predictions on PixMMVP using LLaVA 1.5 (7B) in Fig.
+7.
+Our automatic selection goes through an iterative process of prompting the selected MLLM, in our case GPT-5.1, with N images highlighted with the predicted segmentation to select the best within each group.
+In the final stage, the best images are used to prompt the MLLM to select the final mask that best describes the object of interest.
+keyboard (annotated by the red box)
+bottle
+kite
+drinking glass
+Evaluation.
+We also provide the details on computing the visual question answering accuracy using GPT-4o in the first probing [24] (i.e., ), or when evaluating the language prompt sensitivity in the third probing with GPT-4o (i.e., w/ GPT).
+We use the following prompt: “Given the following question QUESTION, the correct answer is ANSWER.
+Does the following answer correctly answer the question, answer: RESPONSE?
+Respond with a Yes/No”.
+As for the mean intersection over union on PixMMVP (i.e., ), we provide two results: the results on all images, including the ones that do not have an object of interest and their referring expression is marked as “None”, and when excluding this set of images.
+The remaining results in the supplementary for the mean intersection over union are excluding this set unless stated otherwise.
+Note that our benchmarking was conducted on an A6000 48G GPU-equipped machine.
+Prompt sensitivity.
+For the language prompt sensitivity, we generate eight random locations in both the question and the choices to change, excluding the enumeration of the choices.
+Similarly, we randomly select eight locations in the instruction prompting the model to generate an option’s letter.
+Table 5 and Table 6 show the prompt templates that are used in the language prompt sensitivity evaluation in the VQA and the visual grounding, respectively.
+| # | Prompt Template |
+[…]
+[…]
+| 3 | QUES CHOICES INSTRUCT |
+| 4 | Question-QUES CHOICES. INSTRUCT. Answer- |
+[…]
+| 6 | Question, QUES CHOICES. INSTRUCT. Answer, |
+[…]
+[…]
+| 9 | QUES CHOICES INSTRUCT |
+[…]
+| # | Prompt Template |
+| 1 | Can you please segment/detect EXPR in the given image |
+| 2 | Can you segment/detect EXPR in this image? |
+| 3 | Can you identify EXPR in this image? (with grounding) |
+| 4 | Identify EXPR in the scene, with grounding. |
+| 5 | Locate the EXPR and output a tight mask/box. If the object does not exist in the image don’t generate any masks/boxes. |
+| 6 | Output mask/box for the EXPR |
+Appendix B Additional quantitative analysis
+B.1 PixFoundation using open-source models
+In our automatic baseline, we replace GPT-5.1, which is a closed-source model, with another open-source model, in our case, Cambrian (8B).
+Table 7 shows the results on PixMMVP for PixFoundation, the automatic baseline that still surpasses the best pixel-level MLLM, OMG-LLaVA, that was before the release of our benchmark.
+More importantly, this baseline confirms that even with the use of a self-contained model such as Cambrian, without additional help from GPT-5.1, it can still compete with these unified pixel-level supervised models.
+| Method | PixMMVP | ||||
+| OMG LLaVA (7B) | 12.0 | 12.0 | 13.8 | 38.8 | 18.3 |
+| LLaVA 1.5 (7B) + (a+s) | 27.3 | 28.0 | 7.8 | 14.1 | 18.8 |
+| LLaVA 1.5 (13B) + (a+s) | 39.3 | 30 | 8.5 | 13.9 | 20.5 |
+| Cambrian (8B) + (a+s) | 52.0 | 52.0 | 15.5 | 13.3 | 23.9 |
+| Cambrian (8B) + PixFoundation (Ours) | 52.0 | 52.0 | 44.7 | 41.4 | 48.1 |
+| Cambrian (8B) + PixFoundation (Ours) | 52.0 | 52.0 | 24.6 | 25.8 | 34.5 |
+B.2 Random baseline
+Our baselines rely on the maximum attention per output noun phrase to prompt SAM for the segmentation mask.
+Nonetheless, as a lower bound analysis, we evaluate the performance if we use a random point as a prompt instead.
 For fair comparison, we generate random points with the count of output masks that the oracle has to select among (i.e., the number of the output noun phrases).
 We conduct this ablation on PixMMVP using LLaVA 1.5 (7B) base MLLM, with random point prompts followed by the oracle selection among their SAM masks.
-Figure 9 shows that this random + oracle baseline is a strong baseline, yet it lags behind the correct one using the maximum point (i.e., First) with around 12%.
-More importantly, we confirm the stability of the results if we select the second best or third best attention (i.e., Second and Third), which are on-par to the maximum point.
-Thus, even with the oracle selection, using the wrong point prompts lags with a considerable margin behind using the correct ones.
-Summary.
-In summary, we found that highest frequent emergence of grounding is in the last 40-60% of the output text on PixMMVP which might indicate that the reasoning of the model on this challenging task to provide a response impacts when the correct visual grounding can occur.
-More importantly, we show that grounding in MLLMs can emerge in the noun phrase that corresponds to color, position or other characteristics of the object of interest and not necessarily the exact referring expression.
-5 Conclusion
-We proposed two benchmarks showing that pixel-level MLLMs degraded the ability in VQA and even grounding of fine-grained objects.
-Thus, our results questioned whether we are heading in the right direction with these models.
-Additionally, we provide powerful baselines with improved scores without training for pixel-level grounding.
-Impact Statement
-Multi-modal large language models are widely used in various applications, such as robotics, medical image processing and remote sensing.
-The pixel-level understanding within such MLLMs is necessary for such applications that require the localization and even in certain scenarios the delineation of the boundaries for the objects of interest.
-It is even more important to maintain a good chat performance and visual question answering ability in such applications as well.
-In our work, we have investigated the shortcomings of pixel-level MLLMs while providing more challenging benchmarks for these, to improve them further.
-However, as with many other AI advancements there are risks that could be entailed from the deployment of such models.
-There could be inherent biases emerging in such pixel-level MLLMs impacting various under-represented groups.
-We think that our benchmarking efforts and providing a tool to understand the pitfalls in the understanding and reasoning of these models could be an initial direction for mitigating such biases.
-Nonetheless, we leave it for future work to explore this further.
-Appendix A Additional implementation details
-In this section, we cover additional details about our proposed datasets and the implementation of the evaluation setup and baselines.
-We also refer to the output from the questions of the three probing techniques in the supplementary material for all the studied models.
-Datasets.
-Our proposed datasets, PixMMVP and PixCV-Bench, are composed of ground-truth referring expressions describing the object of interest in the respective question and its segmentation mask.
-We show in Fig.
-10 examples of these ground-truth annotations for both datasets.
-It shows the challenging scenarios in pixel-level visual grounding which is strongly tied to the visual question answering task, since an integral part of answering these questions requires the grounding of the object/s of interest.
-Models.
-We also detail the model checkpoints we use for the four pixel-level MLLMs and their variants, retrieved from HuggingFace [22] in Table 2.
-These also include the model checkpoints used for the base MLLMs that were not trained with pixel-level visual grounding.
-It is worth noting, that for GLAMM we use two variants (FullScope and RegCap) since their base model (i.e., FullScope) has low performance in the visual question answering task.
-As such we use the other variant for GLAMM that was fine-tuned for region-level captioning using RefCOCOg dataset.
-Furthermore, we provide details on the oracle selection mechanism, we discard the cases where the ground-truth segmentation is all background in the when analysis, since there is no ground-truth grounding to evaluate against.
-These occur in few cases in PixMMVP.
-Additionally, we provide details on the SAM model that is used in the three baselines and upper bounds in our benchmarks, where we use the ViT-H variant.
-Finally, we provide an illustrative example of our automatic selection mechanism with the corresponding predictions on PixMMVP using Llava 1.5 (7B) in Fig.
-11.
-Our automatic selection goes through an iterative process of prompting the selected MLLM, in our case GPT-4o, with groups of images highlighting the predicted segmentation to select the best within each group.
-In the final stage the best images are used to prompt the MLLM to select the final selected mask that best describes the object of interest.
-| Model Name | Model Checkpoint |
-| LISA | xinlai/LISA-7B-v1-explanatory |
-| GLAMM | MBZUAI/GLaMM-FullScope |
-| GLAMM-RegCap | MBZUAI/GLaMM-RegCap-RefCOCOg |
-| Llava-G | Haozhangcx/llava_grounding_gd_vp |
-| Llava 1.5 (7B) | liuhaotian/llava-v1.5-7b |
-| Llava 1.5 (13B) | liuhaotian/llava-v1.5-13b |
-| Cambrian-1 (8B) | nyu-visionx/cambrian-8b |
-Evaluation.
-We also provide the details on computing the visual question answering accuracy using GPT-4o in the first protocol [19].
-We use the following prompt: “Given the following question QUESTION, the correct answer is ANSWER.
-Does the following answer correctly answers the question, answer: RESPONSE?
-Respond with a Yes/No”.
-Appendix B Additional qualitative analysis
+Figure 9 shows that random + oracle lags behind the correct one using the maximum point (i.e., 1st) with around 28.9%.
+More importantly, we confirm the stability of the results if we select the second-best or third maximum attention (i.e., 2nd or 3rd), which are on par with the maximum attention (i.e., 1st).
+B.3 Conventional visual grounding benchmarks
+We evaluate our upper bound on the conventional benchmarks for RefCOCO/RefCOCO+/RefCOCOg [10, 17] with respect to state-of-the-art methods in pixel-level MLLMs prior to the release of our benchmark.
+Table 8 shows that our upper bound that relies on extracting the output noun phrases, mining their attention maps and selecting the best mask output in relation to these attention maps is competitive to state-of-the-art methods.
+However, the main goal of PixFoundation is to inspect when visual grounding occurs in MLLMs with respect to the output token.
+We still show the results on refCOCO benchmarks for reference.
+Note that unlike refCOCO, our benchmarks pose a greater challenge for pixel-level MLLMs since they are focused on vision-centric tasks that require grounding to answer the questions.
+Furthermore, our paired evaluation enables the evaluation of the grounding in relation to the VQA task to better study these MLLMs.
+Finally, unlike refCOCO variants, our PixMMVP especially provides an out-of-distribution challenging evaluation benchmark, where the majority of these MLLMs have been trained on refCOCO training data.
+| Method | refCOCO | refCOCO+ | refCOCOg | |||||
+| val | testA | testB | val | testA | testB | val | test | |
+| LISA | 74.1 | 76.5 | 71.1 | 62.4 | 67.4 | 56.5 | 66.4 | 68.5 |
+| LISA (ft) | 74.9 | 79.1 | 72.3 | 65.1 | 70.8 | 58.1 | 67.9 | 70.6 |
+| OMG LLaVA (7B) | 75.6 | 77.7 | 71.2 | 65.6 | 69.7 | 58.9 | 70.7 | 70.2 |
+| OMG LLaVA (7B) (ft) | 78.0 | 80.3 | 74.1 | 69.1 | 73.1 | 63.0 | 72.9 | 72.9 |
+| Cambrian (8B) + PixFoundation | 77.8 | 81.1 | 72.8 | 73.2 | 79.5 | 70.1 | 75.5 | 77.5 |
+B.4 When grounding emerges - PixCV-Bench
+In Fig.
+10(a), we show the analysis of when grounding emerges on PixCV-Bench in terms of the location in the output.
+It is worth noting that PixMMVP is more challenging than PixCV-Bench, evidently from the reported IoU and accuracy metrics on both with respect to Table 1 and Table 2 main submission.
+It seems on the less challenging dataset PixCV-Bench, grounding tends to emerge frequently near the beginning of the output.
+This might relate to PixMMVP being more challenging in terms of the level of reasoning than PixCV-Bench or the fact that PixMMVP poses a harder referring segmentation task than PixCV-Bench, which mostly uses the class names.
+Another difference is that PixMMVP is out of the distribution of the seen datasets for most of these MLLMs.
+However, the consistent finding among both datasets is that grounding can emerge coinciding with various concept categories, whether location, color or state, as shown in Fig.
+10(b).
+Note that across this analysis, we compute the frequency per object in the referred expression corresponding to the visual question.
+Hence, if we have two objects in one visual question, such as in the relative positioning questions, each object’s concept, corresponding to the emergence, is computed as part of our analysis.
+Attend and Segment [3] (Cambrian (8B))
+PixFoundation (Cambrian (8B))
+PixFoundation (Cambrian (8B))
+| Model Name | Prompt w/o Expr. | Referring Expression |
+| RGA (7B) | 52.6 | 52.7 |
+| Qwen2.5-VL (7B) | 15.6 | 28.2 |
+B.5 Grounding prompt sensitivity
+We evaluate the language prompt sensitivity in the visual grounding task, aside from the VQA task provided in the main submission.
+Table 9 compares RGA with its baseline MLLM, Qwen2.5-VL, on PixMMVP using the mean intersection over union.
+We conduct two variations per prompt template from Table 6, resulting in 12 variations for the prompt without the referring expression being impacted.
+The language variations randomly modify eight characters in the grounding prompt, where modifications can be either omission, addition or transposition.
+We also perform separate variations on the referred expression only within the aforementioned three modifications.
+Across both prompt variations, whether on the grounding prompt without the referred expression or modifying the referring expression only, RGA shows stronger results and better stability than Qwen2.5-VL.
+The latter exhibits higher sensitivity to the language prompt.
+It is mainly impacted by the prompt template, where it performs better when instructed with the “Locate” keyword than other words in the template to find the object, aside from its sensitivity to spelling errors.
+Appendix C Additional qualitative analysis
 In this section, we provide a qualitative ablation of our baselines and a visualization of the attention maps that can show how vanilla MLLMs are reasoning on the question they are answering.
@@
 Finally, we provide more examples on PixMMVP and PixCV-Bench benchmarks.
-B.1 Baselines ablation
-We show the qualitative ablation among the two baselines and upper bound using the best base MLLM Cambrian-1 (8B) in Fig.
+C.1 Baselines ablation
+We show the qualitative ablation among the baseline and our interpretability mechanism using the best base MLLM Cambrian (8B) in Fig.
 11 on PixMMVP.
-The three confirm that there is grounding emerging in MLLMs that were not trained with pixel-level grounding supervision.
+The three confirm that there is pixel-level grounding emerging in MLLMs that were not trained with mask supervision.
 Nonetheless, it shows that identifying when that grounding emerges is equally important in retrieving the best segmentation of the referring expression.
-The first baseline, attend and segment, assumes the alignment between the attention map that can be mined for the segmentation mask and the noun phrase that has the highest correspondence to the ground-truth category or noun phrase.
+The first baseline, attend and segment, assumes the alignment between the attention map that can be mined for the segmentation mask and the noun phrase that has the highest correspondence to the queried category or noun phrase.
 Our findings quantitatively and qualitatively show otherwise, where grounding can emerge in different output tokens.
-It also shows the oracle upper bound for mask selection, PixFoundation, exhibiting better segmentation than the attend and segment, confirming on the aforementioned finding.
+It also shows the oracle upper bound for mask selection, PixFoundation, exhibiting better segmentation than the attend and segment, confirming the aforementioned finding.
 Additionally, it shows that our simple automatic mechanism, PixFoundation, surpasses the attend and segment as well on PixMMVP.
-Attend and Segment [2] (Cambrian-1 (8B))
-PixFoundation (Cambrian-1 (8B))
-PixFoundation (Cambrian-1 (8B))
-[5pt]
 the yellow skin
-[5pt]
 blue overalls
-[5pt]
 goggles
-[5pt]
 six legs
-[5pt]
 the front legs
-[5pt]
 a hairy texture
-B.2 Attention maps visualization
+C.2 Attention maps visualization
 In this section, we visualize the normalized attention maps, , in Fig.
-13.
-We show two examples for Cambrian-1 (8B) from PixMMVP using the first probing where we directly prompt the model with question and options.
+12.
+We show two examples for Cambrian (8B) from PixMMVP using the first probing where we directly prompt the model with question and options.
 The first row shows outstanding ability to visually ground the different noun phrases from the output text.
@@
 Finally, we find that these attention maps in both examples are not sufficiently accurate to be used for segmentation directly, yet when paired with a powerful segmentation method like SAM it provides a good segmentation performance.
-B.3 When does grounding emerge?
-We show additional examples on when grounding emerges in multi-modal large language models, specifically in the Llava 1.5 (7B) variant, using the second probing to prompt the model to segment what is in the referring expression.
-Figures 14, 16, 15 and 17 show the corresponding predicted masks for the grounding that emerged highlighted in red with the maximum attention point as a black dot.
-Figure 8 in the main text shows the aforementioned four examples with the input prompt, the concept category and the noun phrase corresponding to the best grounding using the oracle selection and the full output text.
-It clearly shows that the correct output token can correspond to location or color but not necessarily the ground-truth referring expression.
-While some of the noun phrases and their masks from the point prompting of SAM do correspond to what the noun phrase is describing.
-It is not always the case, example in Fig.
-16 “the flame” was not able to highlight the correct object yet it appeared in the noun phrase corresponding to the location “the top”.
-It also shows that the majority of correct grounding occurs near the end of the output text after the model reasoned for an answer to the question.
-Additionally, this type of mining for attention maps can show potential mistakes that MLLMs fall into while performing the grounding to reason on what is in the image.
-An example on that is in Fig.
-17, where the model potentially mistakes a crack in the wood for the minute hand.
-Thus, our PixFoundation serves as an interesting tool to interpret and understand how MLLMs work and reason to produce the final output with the oracle selection as an upper bound.
-Yet, there is an open question on how to identify the real mistake from the MLLM vs.
-simple misalignment between the noun phrase and the grounding emerging, since grounding could emerge with phrases describing it or parts of it instead of the object’s noun phrase itself.
-We leave this for future work.
 the image
@@
 a black nose
+C.3 When does grounding emerge?
+We show additional examples of when grounding emerges in multi-modal large language models, specifically in the LLaVA 1.5 (7B) variant, using the second probing to prompt the model to segment what is in the referring expression.
+Figures 13, 14, 15 and 16 show the corresponding predicted masks for the grounding that emerged, highlighted in red with the maximum attention point as a black circle.
+Figure 4 main submission shows the aforementioned four examples with the referred expression, the concept category and the noun phrase corresponding to the best grounding using the oracle selection and the full output text.
+It clearly shows that the correct output token can correspond to location or color, but not necessarily the queried referring expression.
+While some of the noun phrases and their masks, from the SAM point prompting, correspond to what the noun phrase is describing.
+It is not always the case, for example, in Fig.
+15 “the flame” was not able to highlight the correct object, yet it appeared in the noun phrase corresponding to the location “the top”.
+While few scenarios might have the grounding coinciding with multiple noun phrases, such as in Fig.
+13, “a butterfly” and “orange wings”.
+Nonetheless, it is still an important insight that the segmentation can emerge corresponding to noun phrases that do not correspond to the exact referred expression.
+Our PixFoundation serves as an interesting tool to interpret and understand how MLLMs work and reason to produce the final output with the oracle selection as an upper bound.
+In summary, we provide four evidences that pixel-level grounding can emerge corresponding to noun phrases that do not match the exact referred expression, as follows: (i) The attend and segment that rely on SpaCy embeddings lag behind our automatic and oracle mask selection, indicating that the noun phrases closest to the referred expressions are not necessarily where the optimal segmentation emerges.
+(ii) The point accuracy evaluation has confirmed this further, which removes any confounding factors from using three masks per noun phrase in the mask selection due to the point prompt ambiguity.
+The point accuracy only evaluates the emergent grounding with respect to the noun phrases.
+(iii) We show quantitative analysis on the location and the concept categories of the noun phrases where the grounding emerges, which confirms the previous result.
+(iv) We show qualitative analysis to confirm this further.
+C.4 PixMMVP benchmark
+Figure 17 shows additional results on PixMMVP benchmark comparing different pixel-level MLLMs with our oracle baseline using LLaVA 1.5 (7B).
+While GLAMM shows strong pixel-level visual grounding yet we have shown earlier that it is almost incapable of visual question answering, which renders the model weak for general-purpose tasks.
 The flame
@@
 the 12 o’clock position
-B.4 PixMMVP benchmark
-Figure 18 shows additional results on PixMMVP benchmark comparing different pixel-level MLLMs with our oracle baseline using Llava 1.5 (7B).
-While GLAMM shows strong pixel-level visual grounding yet we have shown earlier that it is almost incapable of visual question answering which renders the model weak for general purpose tasks.
 On the other hand, OMG-LLaVA shows a better balance in pixel-level visual grounding and visual question answering as previously detailed.
@@
 In fact, certain MLLMs that were trained with pixel-level visual grounding, such as LISA, have degraded the performance with respect to the hidden information already existing in powerful MLLMs that were not trained with such supervision.
+C.5 PixCV-Bench benchmark
+Figure 18 shows qualitative results on PixCV-Bench.
+It shows that pixel-level MLLMs struggle with segmenting the object annotated by the red box, unlike our oracle baseline, PixFoundation.
+Indeed the attention maps from these MLLMs are looking at the right object annotated by the red box without receiving any pixel-level grounding supervision during training.
+Appendix D Analysis on the output length
+In this section, we provide additional analysis on the output length on average through PixMMVP dataset using the first and second probing schemes.
+Specifically, we report the output length as the number of characters in the output, and the number of noun phrases extracted from it.
+The reason to study this, since it has relation to the number of noun phrases and consequently the number of masks our interpretability mechanism is selecting among.
+Table 10 shows the average output length computed across PixMMVP dataset, comparing the three base MLLMs.
+We notice that Cambrian (8B) generates longer outputs with a considerable margin than LLaVA variants.
+Hence, we believe the superiority of the oracle upper bound with Cambrian in the grounding has strong correlation to producing longer outputs with more attention maps to mine and select from, than LLaVA variants.
+Nonetheless, it makes it more challenging for the automatic baseline.
+| Model Name | Probing | Output Length | # Noun Phrases |
+| LLaVA 1.5 (7B) | First | 44.2 | 2.3 |
+| LLaVA 1.5 (13B) | First | 45.3 | 2.4 |
+| Cambrian (8B) | First | 313.8 | 15.2 |
+| LLaVA 1.5 (7B) | Second | 92.6 | 5.2 |
+| LLaVA 1.5 (13B) | Second | 97.2 | 5.5 |
+| Cambrian (8B) | Second | 561.3 | 27.3 |
 the butterfly’s feet
@@
 the yellow animal’s head
-B.5 PixCV-Bench benchmark
-Figure 19 shows qualitative results on PixCV-Bench.
-It shows that pixel-level MLLMs struggle with segmenting the object annotated by the red box unlike our oracle baseline, PixFoundation.
-Indeed the attention maps from these MLLMs are looking at the right object annotated by the red box without receiving any pixel-level grounding supervision during training.
 cell phone
@@
 chair and cell phone
-Appendix C Failure case analysis
-In this section, we conduct additional failure case analysis of pixel-level MLLMs and our baselines qualitatively and quantitatively.
-C.1 Failures in visual question answering
+Appendix E Failure Cases Analysis
+In this section, we conduct an additional failure case analysis of pixel-level MLLMs and our baselines qualitatively and quantitatively.
+We start with evaluating whether the failures of these MLLMs occur in visual grounding, VQA or both.
+Since our evaluation is paired, it enables such an analysis to have a better understanding of where the failures are stemming from.
+Figure 19 shows the frequency of failures per category, where the majority stem from failures in both, especially in the pixel-level MLLMs, except RGA, which shows the majority of failures are in the VQA.
+The standard MLLMs perform better in the VQA than grounding, as expected.
+However, improving pixel-level visual grounding in a naive manner that neglects the language capability eventually resulted in failures in both.
+On the other hand, RGA shows the highest success on both VQA and visual grounding, showing it as a promising direction, yet, as shown in the prompt sensitivity results, there are still aspects that can be improved to ensure they do not lose the language capabilities and robustness that existed in standard MLLMs.
+We encourage better training recipes, datasets and design choices to enable visual grounding in MLLMs without sacrificing their language abilities that are tightly coupled with grounding.
+E.1 Failures in Visual Question Answering
 We start with a fine-grained quantitative analysis of how the studied models perform across PixMMVP and PixCV-Bench.
@@
 It also shows that ADE20K dataset serves as a more challenging dataset than COCO.
-C.2 Failures in pixel-level visual grounding
+E.2 Failures in Pixel-level Visual Grounding
 Finally, we show qualitatively the failure cases of the oracle upper bound in Fig.
@@
 Hence, the oracle upper bound is generally inheriting SAM failures.
-However, its main purpose of showing that the hidden information within powerful MLLMs is sufficient to perform pixel-level grounding is achieved, and even surpassing pixel-level MLLMs without degrading their VQA abilities.
-Appendix D Additional quantitative analysis
-D.1 When grounding emerges - CV-Bench
-In Fig.
-23(a) we show the analysis on when grounding emerges on PixCV-Bench in terms of the frequency of the grounding location.
-It is worth noting that PixMMVP is more challenging than PixCV-Bench evidently from the reported IoU and accuracy metrics on both reported in Table 1.
-It seems on the less challenging dataset PixCV-Bench, grounding tends to emerge frequently near the beginning of the output.
-This might relate to PixMMVP being more challenging in terms of the level of reasoning than CV-Bench or the fact that PixMMVP poses a harder referring segmentation task than PixCV-Bench, which is mostly using the class names.
-However, the consistent finding among both datasets is that grounding can emerge coinciding to various concept categories whether location, color or state as shown in Fig.
-23(b).
-D.2 Analysis on the output length
-In this section, we provide additional analysis on the output length on average through PixMMVP dataset using the first and second probing schemes.
-Specifically, we report the output length as the number of characters in the output, and the number of noun phrases extracted from it.
-The reason to study this, since it has relation to the number of noun phrases and consequently the number of masks our baselines are selecting among.
-Table 3 shows the average output length computed across PixMMVP dataset, comparing the three base MLLMs.
-We notice that Cambrian-1 (8B) generates longer outputs with a considerable margin than Llava variants.
-Hence, we believe the superiority of the oracle upper bound with Cambrian-1 in the grounding has strong correlation to producing longer outputs with more attention maps to mine and select from, than Llava variants.
-| Model Name | Probing | Output Length | # Noun Phrases |
-| Llava 1.5 (7B) | First | 44.2 | 2.3 |
-| Llava 1.5 (13B) | First | 45.3 | 2.4 |
-| Cambrian-1 (8B) | First | 313.8 | 15.2 |
-| Llava 1.5 (7B) | Second | 92.6 | 5.2 |
-| Llava 1.5 (13B) | Second | 97.2 | 5.5 |
-| Cambrian-1 (8B) | Second | 561.3 | 27.3 |
+However, its main purpose of showing that the hidden information within powerful MLLMs is sufficient to perform pixel-level grounding is achieved, and even surpasses some of the pixel-level MLLMs without degrading their VQA abilities.
+Appendix F Impact Statement
+Multi-modal large language models are widely used in various applications, such as robotics, medical image processing and remote sensing.
+The pixel-level understanding within such MLLMs is necessary for such applications that require the localization and even in certain scenarios the delineation of the boundaries for the objects of interest.
+It is even more important to maintain a good chat performance and visual question answering ability in such applications as well.
+In our work, we have investigated the shortcomings of pixel-level MLLMs while providing more challenging benchmarks for these, to improve them further.
+However, as with many other AI advancements there are risks that could be entailed from the deployment of such models.
+There could be inherent biases emerging in such pixel-level MLLMs impacting various under-represented groups.
+We think that our benchmarking efforts and providing a tool to understand the pitfalls in the understanding and reasoning of these models could be an initial direction for mitigating such biases.
+Nonetheless, we leave it for future work to explore this further.
+Appendix G Limitations
+Note that our baselines do entail a computational overhead with the use of the mask selection process.
+Nonetheless, the benefit from exploring what is already learned in these MLLMs through mining the attention maps with an understanding of when grounding emerges, provides greater benefit to interpretability.
+We believe interpretability of MLLMs is a crucial aspect when following a responsible approach to AI.